A Complete Bibliography of Publications in Canadian Journal of Mathematics = Journal canadien de

mathématiques for the decade 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

02 June 2023 Version 1.31

Title word cross-reference

(4, 4) [108]. 1 [524, 483]. 1 + ϵ [75]. 2 [398, 57, 66, 562, 193]. 3 [248, 490, 222, 256, 324]. 4 [69]. 55 [96]. 9 [81]. ^{∗} [203, 94]. A(K) [35]. A_{2} ∗ T_{2} [389]. A_{p}^{r}(G) = A_{p}∩L^{r}(G) [77]. A_{R}(K) [35]. b [371]. BZ/(2^{q}) [462]. C(K) [33]. C∗ [455, 466, 468, 316, 566]. C^{∗} [202, 546, 187, 506, 422, 510, 442, 240, 565, 156, 144, 151]. C_{0} [302]. c_{2} [503]. C_{R}(K) [35]. D [178]. ER(2) [462]. G [392]. G_{2} [547, 490]. GL(3) [5]. GL_{2} [544]. GL_{n} [474, 480]. GU(1, s) [36]. GU(n − 1, 1) [292]. H [56, 164]. H^{1} [126, 13]. Â [353]. ∞ [350]. j [81]. K [308, 38, 266, 164, 144, 240]. K^{2} = 6 [193]. K_{1} [360]. L [186, 78, 509, 284, 11, 408, 23, 544, 314, 172, 387, 241]. L^{p} [260, 204, 102, 7, 64]. L^{q} [431]. LMO [13]. M [229]. B^{N} [226]. C^{2} [146]. CP^{n} [462]. H^{2} × R [448]. P^{1} × P^{1} [447, 221]. P^{N} [377]. S^{2} × R [448]. S^{n} × R [259]. Z/2 [281]. Z_{p}^{2} [275]. GL_{n} [500]. L [547]. GL(N) [252]. I [187]. μ = 2 [289]. N [430, 345, 64, 386]. O(2n) [86]. \overline{GSp}_{2n}(F) [306]. \overline{M}_{0,7} [425]. p [345, 464, 5, 265, 338, 516, 376, 388, 86, 56, 113, 114, 23, 544, 172, 500, 436, 46]. p_{g} = q = 2 [193]. P F A [173]. PFA(S)[S] [173]. q [372, 80, 223, 81, 261]. Q_{α}(R^{n}) [59]. S [491, 445]. S^{1} [206, 54]. S^{1} × Σ [392]. S_{3} [111]. SL(2, C) [353]. SL_{2} [66]. SL_{2}(C) [467]. SO(2n) [86]. SO(N) [436]. Sp(N) [436]. Sp_{2}(F) [46]. Sp_{2n} [113]. θ [376]. V = L [173]. w∗ [495]. W_{ω}^{2}, p [136]. G̃L(r) [361]. x + y = z^{2} [526]. Z [2, 138]. ζ′(s)^{2} [560]. ζ′(s)ζ′′(s) [560, 560].

-Adic [113, 345, 5, 265, 338, 376, 388, 86, 114, 544, 172, 500, 436, 23, 46]. -adiques [376]. -algebra [156]. -algebras [202, 565, 94, 566, 203, 144, 308, 455, 546, 187, 506, 466, 422, 510, 151, 442, 468, 240, 316]. -analogue [372, 81]. -body [430]. -cohomology [462]. -colourable [178]. -connectedness [524]. -covers [111]. -Dimensional [66, 386, 64]. -extension [275]. -folds [324]. -function [64, 547]. -functionals [38]. -functions [186, 78, 509, 284, 11, 544, 172, 241]. -Functoriality [408]. -functors [360]. -ideal [491]. -invariant [445]. -Laplacian [56]. -manifolds [206, 69, 54]. -Modules [2, 138]. -Monopoles [392]. -motives [483]. -polynomial [353]. -principle [371]. -régulateurs [376]. -representations [562]. -row [398]. -Schur [261]. -semicircular [223]. -Semicrossed [495]. -Solvability [136]. -space [256]. -spaces [204]. -spheres [248]. -Split [108]. -stacks [345]. -Summable [516]. -Theory [144, 266, 240]. -torsion [222]. -transportation [350]. -Type [56]. -ultraspherical [80]. -values [23, 314, 387]. -Weakly [229].

3-manifolds [49]. 3D [184]. 3D-Regularized [184].

Abelian [354, 514, 218, 414, 249, 378, 322, 305]. abéliennes [514]. Absolute [385, 441]. Absorption [566, 468]. Abstract [458]. Action [395, 525]. Actions [330, 342, 427]. Addendum [206, 272]. Additive [411, 459, 82]. Adic [113, 345, 5, 265, 338, 376, 388, 86, 114, 23, 544, 172, 500, 436, 46]. adiques [376]. adjoint [218]. Admitting [386]. AF [328]. Affine [545, 480, 27]. AH [154]. AH-Algebra [154]. AI [87]. aide [68]. Albanese [193]. Alexandrov [441]. Algebra [289, 156, 102, 403, 385, 329, 88, 537, 154, 261]. Algebraic [2, 138, 489, 321, 475, 520, 477, 28, 6, 376]. Algebraically [536]. Algebras [299, 185, 203, 319, 228, 92, 398, 490, 328, 458, 402, 77, 388, 404, 207, 313, 251, 415, 536, 66, 417, 278, 277, 150, 22, 418, 550, 125, 144, 114, 147, 531, 261, 202, 308, 455, 546, 565, 187, 506, 466, 94, 422, 87, 510, 10, 151, 442, 468, 240, 43, 463, 316, 566]. algébrique [376]. Algebroids [106]. Almost [447, 420, 426]. Along [492, 220]. Amenability [440, 43]. Amenable [229]. Ampère [20, 399, 14]. Ample [91]. ampleness [60]. Analogue [556, 372, 81]. Analogues [406]. Analysis [155, 168]. Analytic [215, 410, 179, 497]. Anderson [548]. Angelesco [394]. Anisotropic [446, 435]. Annihilators [557]. Anticyclotomic [152]. aposyndetic [12]. Application [236, 454, 21]. Applications [439, 74, 412, 419, 285, 451, 196, 455, 320, 40, 10, 58]. Approach [504, 538, 6]. Approximants [394]. Approximate [270, 560, 517]. Approximating [318]. Approximation [38, 19, 358]. approximations [71]. Arbitrary [349, 488]. Archimedean [406, 116]. Area [481]. Arens [531]. Arising [202, 364, 461]. Arithmetic [108, 535, 406, 538, 364]. Arrangements [110, 520]. Arthur [68, 436]. Artin [345, 236]. Aspects [365]. Associated [289, 422, 292, 98, 511, 520, 295, 343, 83, 79, 567]. associative [228]. Asymptotic [73, 432, 550, 3]. Asymptotical [255]. Asymptotically [449]. Asymptotics [79, 98, 394]. Attaining [506]. Attraction [346]. Automatic [228]. Automorphic [284, 361]. Automorphism [354]. Automorphismes [128]. Automorphisms [308]. AV [106]. averages [38]. Axiom [519].

Back [513, 522, 533, 542, 552, 564]. Balayage [161]. ball [64]. Banach [228, 93, 77, 329, 386]. Barriers [117]. Barycenters [441]. Bases [264, 480]. Basis [545, 119, 555]. Batalin [537]. Baumslag [137]. Be [420]. behavior [3]. behaviour [255]. Bellman [451]. Beltrami [237]. Bending [341]. Bergman [397, 24]. Besov [237]. Bessel [72]. Best [93]. Betti [281]. Between [402, 265, 146]. Beurling [472, 402]. Bicommutant [495]. bidisk [24]. Biflatness [43]. Bilinear [368]. Billiards [242]. Binary [101, 241]. Bipositive [402]. Bishop [456, 558, 257]. Blanchfield [488]. Bloch [253, 104, 432]. Bloch-type [253]. BMO [42, 136, 71]. BMO-estimates [71]. BMW [261]. Bochner [329]. Bodies [297, 481, 169, 70]. body [252, 430]. Bollobás [456, 558, 257]. Boltzmann [504]. Bombieri [167]. Borel [122]. Bott [122]. Bound [356]. Boundary [510, 226, 465, 63, 64]. Bounded [298, 222]. Boundedness [162]. Bounding [420]. Bounds [377, 499, 49]. Boussinesq [184]. Bowen [180]. Brackets [166]. Bragg [213]. Branch [279]. Branched [263]. Branching [5]. Brascamp [44]. Bratteli [336, 246]. Brauer [545]. Brownian [158]. Bruhat [119]. Brunn [439]. Brylinski [155]. Bulk [209]. Bundles [335, 281, 373, 530, 60].

C [203, 94]. Calabi [505, 96]. Calculus [562, 27]. Calderón [162]. Can [420]. Cancellation [543]. Canonical [479, 480, 65, 50]. Cantor [310]. Capacity [435]. caractère [252]. Cardinal [147]. Carleman [357]. Carleson [42, 295]. carpet [79]. Cartan [369]. Case [503, 85, 25]. Cases [535, 444]. Casimirs [166]. Casselman [119, 555]. Cassels [338]. Casson [353]. Categorical [365]. Categories [298, 427]. Categorification [261]. Category [336, 366]. Cauchy [136]. central [23, 174]. Centre [390].

Centre-valued [390]. Certain [377, 356, 329, 442]. Cesàro [523]. Character [55, 252, 467]. Characterisation [21]. Characteristic [356]. Characterization [445, 270]. Characterizations [80, 70, 343]. Characterizing [385]. Characters [186, 88]. Chern [344, 409]. Chiral [120]. Chowla [123, 26]. Christoffel [209]. Circulant [517]. CJM [512, 513, 521, 522, 532, 533, 541, 542, 551, 552, 563, 564]. Class [478, 299, 557, 236, 302, 458, 231, 538, 486, 460, 11, 24]. Classes [288, 344, 357, 352, 491, 379]. Classic [171]. Classical [430, 385]. Classification [438, 156, 19, 24, 466, 87]. Classifying [267, 537]. Claw [507]. Claw-free [507]. Clifford [545]. Closed [129, 450, 536, 442, 286]. Closures [175]. CM [471, 85, 322, 11]. CM-case [85]. Co [497]. Co-analytic [497]. Coarse [340]. Cobordism [321]. Cocompact [256]. Cocycle [132]. Coefficient [237, 252]. Coefficients [559, 97, 183, 136, 13]. Coherent [549]. Cohomological [388, 538]. Cohomology [17, 34, 530, 201, 477, 62, 462, 433]. Colmez [26]. color [562]. colourable [178]. Coloured [199]. Colouring [507]. colourings [21]. Combinations [104, 314]. Combings [303]. Commensurability [304]. Commutative [102, 264, 28, 348]. Commutativity [39, 45]. Commutators [139, 384]. Compact [440, 458, 231, 417, 391, 486, 316, 456, 457, 14]. Compactifications [340]. Compactness [139]. Comparison [455, 415, 243]. complement [465]. Complementarity [286]. Complemented [116]. Complete [447, 3, 87]. Completely [531]. Completeness [508, 212]. Completion [503]. Complex [215, 100, 245, 189, 224, 305, 14]. Complex-valued [100]. Complexes [528, 222, 475, 89]. Component [400]. Components [553, 406]. Composition [490, 104, 179, 258]. Compositional [159]. Computable [450]. Computation [488]. Computer [227]. Computing [29]. concave [419]. Concentration [439]. Condition [339]. conditionally [82]. Conductor [405]. Cone [91]. Conformal [196]. Conformal-harmonic [196]. conforme [196]. conforme-harmoniques [196]. Congruence [185, 292]. Congruences [121, 491]. Conjecture [186, 535, 556, 52, 236, 123, 159, 327, 371, 501, 75, 85, 26]. Conjectures [376, 200]. Conjugacy [288]. Connected [472, 331]. Connectedness [238, 524]. connections [51]. Connectivity [293]. Consequences [173, 62]. Consistent [254, 173]. Constant [437, 130, 141, 469]. constituents [8]. Constrained [358]. Constructions [202, 120, 67]. Contact [210]. Contained [187]. containing [486]. Context [430]. continua [12]. Continuity [247, 228, 385, 441]. Continuous [487, 92, 229, 550, 80]. Contraction [473]. Contractions [302]. contragrédientes [291]. contragredients [291]. Controlled [90]. Convergence [559, 184, 350, 540]. Convex [238, 481, 274, 442, 169, 300, 42, 70]. convexe [300]. Convexity [297, 257, 48]. Convolution [412, 458, 384, 385, 231, 417, 364, 247, 82]. Corona [156]. corps [252]. Corrádi [325]. Corrádi-Hajnal [325]. Correction [202]. Correspondence [546, 565, 388, 95, 405, 408, 171, 8]. Corrigendum [138, 565, 278, 567]. Costs [311]. Countable [272, 271]. Counted [554]. Coupled [177]. Courant [106]. Covariant [440]. Covariants [244]. Cover [512, 513, 521, 522, 532, 533, 541, 542, 551, 552, 563, 564]. Covers [263, 111]. Coxeter [515, 255, 256]. CR [556, 106]. Creation [403]. Cremona [333]. Criteria [60].

Critical [110, 285, 4, 58]. critically [377]. Crossed [546, 565, 154]. Cubature [181]. Cubic [554, 142, 125, 241, 73, 21, 40]. Cumulants [223]. Cuntz [455]. Curvature [437, 49, 130, 301, 188, 378, 226, 3, 67, 469]. Curve [281, 381, 323, 23]. Curved [217, 430]. Curves [129, 318, 108, 131, 103, 367, 190, 375, 528, 356, 483, 369, 406, 475, 134, 287, 393, 280, 494, 58]. Cusp [97, 387]. Cuspforms [396]. Cuspidal [370]. Cycles [142, 305, 172, 234]. Cyclic [263, 557]. Cyclotomic [432]. cylindrical [260].

D [398]. Darboux [212]. Darmon

[176, 172]. d’Arthur [68]. Daugavet [198]. Deciding [519]. Decomposable [19, 502]. Decomposition [112, 12]. Defining [289]. Definiteness [391]. definition [14]. Deformations [549, 417, 29]. Degenerate [545, 290, 283, 56, 343, 51]. Degenerating [364]. Degenerations [383, 323]. Degree [222, 193, 225]. Degrees [479, 411, 525]. del [133]. Delaunay [443]. Delone [191]. Dem’janenko [494]. Dendriform [403]. Dendrites [379]. Denjoy [357]. Dense [396, 272, 271]. Densities [165]. Density [325]. d’entrelacement [68]. Derivations [219]. Derivative [274]. Derived [298]. Designs [181]. Determinacy [410]. determinant [10]. d’exponentielles [320]. Diagonal [122, 561, 73]. Diagonals [208]. Diagram [398]. Diagrams [336, 246]. Diffeology [239]. Difference [347, 450, 170]. Different [420, 295]. Differentiability [270]. Differential [518, 235, 7]. Differentials [422]. Differing [411]. Diffraction [213]. diffusion [359]. Digits [424]. Digraphs [178]. Dihedral [200]. dilated [38]. Dimension [215, 312, 150, 452, 242]. Dimensional [189, 168, 66, 210, 331, 443, 242, 157, 311, 506, 64, 508, 536, 407, 98, 386]. Dimensions [553, 94, 482, 334]. Dimensonal [125]. Dirac [125, 233]. Direction [130].

Directions [220]. Dirichlet [478, 421, 161, 41, 285, 260, 183, 136]. Discrete [216, 129, 546, 565, 86, 378, 115, 500, 252, 80, 252]. Disjointness [420, 426]. Disk [317, 41]. disks [63]. Dispersive [285]. Distance [217, 194, 350]. Distinction [476]. distortion [386]. Distribution [1, 349, 367, 509, 32, 114, 16]. Distributive [185]. Ditkin [77]. Divergence [107, 539]. Divergence-free [539]. Divisible [548]. Division [393]. Divisor [349]. Divisors [425, 47]. Domain [472]. Domains [37, 250, 260, 146, 214]. domination [300]. d’opérateurs [233]. Douady [128]. Doubled [493]. doubling [71]. Drifts [539]. Drinfeld [278, 349, 277]. Drops [150]. Dual [402, 403, 153]. Duality [298]. d’un [282, 376]. Dyadic [451]. Dyck [159]. Dynamical [535, 339, 191, 331]. Dynamics [437, 310].

Eberlein [329]. Edge [485, 21]. edge-colourings [21]. Edge-Maximal [485]. Edges [370]. Eigenforms [200]. eigenspaces [7]. eigenvalues [41, 9]. Eigenvarieties [309, 396]. Einstein [274, 296]. Ekedahl [470]. Elasticity [498]. Elementary [349]. Elements [28, 506, 490, 415]. Elliptic [381, 131, 134, 393, 494, 486, 343, 454, 4, 23, 58]. Embedding [338, 293]. Embeddings [450, 386]. Empirical [350]. Endomorphism [154]. Endomorphisms [377]. Ends [250]. Energy [129, 285, 499, 296]. Enriques [511, 567]. entries [454]. Entropy [94]. Environments [283]. epsilon [282]. Equation [237, 459, 170, 334, 540, 461]. Equations [411, 140, 560, 177, 399, 211, 539, 260, 434, 136, 486, 73, 7, 13]. Equicontinuous [191]. equilibria [469]. Equilibrium [347]. Equivalence [362, 302, 550]. equivalences [69]. Equivariant [346, 363, 17]. Erdos [327].

Ergodic [199]. Erratum [16, 206, 566, 433]. Escape [359]. espace [128]. essential [45]. Estimate [285]. Estimates [107, 397, 276, 539, 242, 13, 71]. Étale [126]. Euclidean [175, 105]. Euler [356, 469]. Euler-type [469]. Example [278]. Examples [364]. Exel [313]. Existence [248, 121, 183, 387]. Explicit [488, 338, 393, 225, 494, 40, 67, 81]. Exponent [367, 4]. Exponential [339, 499]. Exponents [446]. Expression [282, 81, 282]. Extended [328]. Extension [557, 266, 275]. Extensions [53, 432, 51]. Extrema [41]. Extremal [451]. extremals [70]. Extreme [207]. Extremely [77].

faced [348]. facteur [282]. Factor [487, 282]. Factorization [37]. Factors [476, 460]. Failures [279]. Faithfulness [330]. Families [186, 348, 190, 388, 164, 332, 182, 127, 380, 58]. Family [423]. Fano [65, 323]. Fast [117]. Fat [447, 208]. Fekete [431]. Fermat [140, 367]. Few [406]. Fiber [395]. Fibered [248]. Fibers [37]. Fibred [225]. Fibres [398]. Field [5, 130, 557, 450, 536, 275, 538, 46]. Fields [381, 92, 338, 421, 249, 450, 141, 168, 276, 284, 66, 423, 499, 287, 460, 40, 11]. Filters [203]. Filtrated [144]. Filtrations [160]. Finite [410, 288, 246, 439, 381, 506, 339, 189, 482, 377, 168, 499, 466, 10]. Finite-dimensional [506]. Finite-time [339]. Finitely [185, 19]. Finsler [6, 67, 243]. First [349]. Fixed [534, 473]. Flag [188, 290, 67, 27]. Flags [407, 171]. Flatness [410]. Fleming [76]. Floer [201]. Flow [326, 74, 516, 561, 409, 79]. Fluctuation [454]. folds [505, 324]. Foliations [245, 248]. fondamental [149]. Forcing [519]. Forest [553]. Form [554, 107, 338]. Formal [37, 366]. Forms [265, 474, 419, 121, 368, 388, 161, 518, 284, 432, 97, 491, 332, 152, 314, 387, 241, 502, 25].

Formula [516, 312, 313, 282, 83, 26]. Formulae [181, 212, 73]. formulars [252]. Formulas [480]. formule [282, 401]. formules [252]. Four [157, 505, 121]. Four-dimensional [157]. Four-folds [505]. Fourier [102, 197, 417, 97, 115, 294, 531]. fractal [59]. Fractional [434, 435, 50]. Free [202, 337, 72, 495, 447, 312, 94, 422, 269, 416, 51, 467, 69, 539, 82, 502, 507]. Freeness [529]. French [196, 282, 128, 514, 300, 252, 376, 291, 68, 233]. Frequency [107, 381]. friables [320]. Friedlander [167]. Frobenius [367, 251]. Front [512, 521, 532, 541, 551, 563]. Full [266, 94, 482]. Fully [486]. Function [458, 421, 141, 352, 416, 276, 423, 98, 451, 343, 64, 547, 40, 59]. Function-theoretic [352]. Functional [148, 560, 274, 7]. Functionals [238, 229, 296, 38, 70]. Functions [307, 471, 264, 535, 556, 1, 253, 489, 351, 268, 403, 88, 209, 258, 164, 306, 247, 186, 16, 83, 78, 79, 31, 20, 509, 284, 11, 42, 544, 172, 241]. Functor [298]. Functoriality [50, 408]. functors [360]. Fundamental [54, 206, 208, 484, 151, 69]. Furstenberg [240].

Galois [565, 50, 546, 84, 287]. Gamma [476]. Gauss [301, 352]. Gaussian [439]. General [39, 559, 221, 356, 14]. Generalised [219]. Generalized [245, 472, 251, 519, 106, 296, 7, 64]. Generating [88]. Generators [289, 147]. Generic [310, 349, 8]. Genericity [113]. Genuine [153, 306]. Genus [108, 230]. Geography [324]. Geometric [466, 405, 370, 304]. Geometry [362, 543, 553, 157, 101, 105, 208, 28, 467]. Ginzburg [383]. Girth [178]. given [9]. Givental [383]. Global [307, 434, 184, 234, 48]. Goldman [341]. good [470]. Gorenstein [324]. Goresky [27]. Gosset [133]. Gradient [220, 486].

Graph [503, 466]. Graphs

[422, 132, 222, 485, 293, 507, 466, 21]. Green [268]. Gromov [323]. Grothendieck [298]. Group [557, 366, 525, 54, 206, 269, 208, 151, 115, 181, 342, 153, 240, 467, 11, 69]. groupe [149]. groupes [401]. Groupoids [34, 433]. Groups [288, 137, 543, 546, 148, 381, 524, 440, 365, 515, 458, 189, 229, 368, 508, 18, 527, 95, 231, 56, 232, 112, 256, 275, 378, 133, 417, 432, 484, 114, 152, 400, 170, 391, 360, 316, 247, 28, 122, 460, 255, 31, 322, 565]. Growth [148, 556, 253, 267, 256, 432, 322].

Haar [254]. Hadamard [262]. Hajnal [325]. Hall [264]. Hamilton [177]. Hamiltonian [335, 347, 235]. Handlebodies [132]. Hankel [25]. Hansen [501]. Hardy [446, 4, 397, 472, 56, 451, 24, 25, 343]. Harmonic [253, 224, 168, 196]. Harmonics [561]. harmoniques [196]. Hasse [479]. Hausdorff [242]. Heaps [223]. Heat [268, 434, 79]. Hecke [278, 186, 88, 66, 277]. Heegner [53, 369]. Height [377]. Heisenberg [527]. help [68]. Hermite [141, 394]. Hermitian [158, 9, 486]. Hessian [220, 511, 567]. Heteroclinic [180]. Higgs [373]. High [553]. Higher [238, 205, 97, 99, 334, 242, 25]. Hilbert [244, 227, 274, 418, 332, 152, 286, 387, 200]. Hodge [263, 470, 364]. Holomorphic [307, 63, 97, 146]. Holomorphie [68]. Holomorphism [68]. Homogeneities [295]. Homogeneous [157, 188, 458, 272, 271, 168, 62, 162, 12]. Homology [248, 222, 493]. Homomorphisms [228, 219, 543]. Homotopy [156, 189, 303, 69]. Hopf [195, 262, 251]. Hulls [442]. Hurwitz [230]. Hyperbolic [301, 256, 315, 49]. Hyperbolicity [250]. Hyperelliptic [356]. Hypergeometric [471]. Hyperkähler [505]. Hyperplane [110]. Hyperplanes [263]. Hyperrigidity [496]. Hyperspace [310]. Hypothesis [519, 279].

Ideal [289, 557, 452, 50, 87, 491]. Ideals [372, 221]. Idempotents [490]. identity [71]. II [429, 440, 92, 85, 482]. III [62, 151]. Ilyashenko [463]. Images [428]. Imaginary [557, 275, 284]. Implicit [177]. Inclusions [566, 468]. Incompressible [375]. Ind-groups [122]. Indecomposable [372, 84]. Indefinite [388, 461]. Index [198, 390]. Induced [179]. Inductive [150]. Inequalities [439, 148, 419, 481, 351, 56, 294]. Inequality [382, 59]. inertia [9]. Infinite [202, 298, 339, 508, 423, 125, 255]. Infinite-dimensional [508]. Infinite-Dimensonal [125]. Infinitely [543, 226]. Infiniteness [154]. Infinitesimal [274]. infinitésimale [401]. Initial [260]. Injectivity [440]. Inner [265, 474, 258]. Integers [2, 138]. Integrable [373]. Integral [516, 384, 393, 282, 45]. intégrale [282]. Integrals [139, 474, 205, 239]. interacting [76]. Interior [13]. Interlacing [143, 68]. Intersection [297, 318]. Intersections [344, 447]. Intertwining [444]. Interval [150]. Intrinsic [365]. Invariance [503]. Invariant [353, 198, 503, 445, 460]. Invariants [479, 205, 132, 112, 416, 303, 370, 323, 145, 147]. Inversion [269]. Invertible [380]. Irrational [328]. Irreducibility [501]. Irreducible [19, 8]. Irregular [324]. Isobaric [309]. Isometric [531, 47]. isometries [47]. Isomorphic [523]. Isomorphism [404]. Isomorphisms [402, 418, 128]. Isoperimetry [439]. Isoresonant [100]. Issue [512, 513, 521, 522, 532, 533, 541, 542, 551, 552, 563, 564]. Iterated [308, 205, 98]. Iwahori [119, 555, 66]. Iwaniec [167]. Iwasawa [200].

Jacobi [177, 358]. Jacobians [108, 400]. Jacquet [388, 401]. Jiang [566, 468]. John [59]. Jonquières [333]. Jordan [302].

Jump [359]. Jump-diffusion [359].

K3 [91, 414, 225]. Kähler [3, 14, 296].

KAM [216]. Kampen [126]. Kantor [428]. Kato [432]. Kazhdan [555]. Kepler [437, 461, 273]. kernel [464]. Kernels [366, 268, 384]. Khovanov [398, 493]. Kirchberg [92]. Kirchhoff [374]. Knots [353]. Knotted [132]. Knuth [171]. Kostant [155]. Kottwitz [252]. Kummer [53].

Lagrange [195]. Lagrangian [63].

Laguerre [135]. Landau [383]. Lang [535]. Langlands [388, 405, 8, 113]. Laplace [211]. Laplacian [56, 41]. Laplacians [98]. Large [178, 117]. Lattices [118]. Laumon [483]. Laurent [383]. Law [223]. Laws [72, 366, 422]. Lech [308]. Lefschetz [211]. Left [508]. Lehmer [85]. Leindler [382]. Lemma [484]. lemme [149]. Lengths [498]. Level [140, 121, 96, 152]. LF [89]. LF-spaces [89]. Lie [299, 247, 363, 148, 319, 524, 64, 508, 18, 385, 269, 31, 22, 125, 114]. Lieb [44]. Lift [264, 547]. Lifting [288, 145]. Limit [174, 220, 528, 475, 82]. Limited [197]. Limits [150]. Line [520]. Linear [534, 382, 458, 528, 104, 356, 475, 183, 314, 83, 539]. Links [445, 488]. Lipschitz [524, 527, 214]. Littlewood [264, 317]. Local [297, 534, 473, 482, 548, 405, 66, 113, 408, 48, 252, 439, 376]. Localized [294]. Locally [440, 217, 84, 231, 417, 391, 457, 33]. locaux [376]. locus [396, 36]. Loewner [334]. Log [419, 351]. Log-concave [419]. Log-Subharmonic [351]. Logarithmes [514]. Logarithmic [351]. Logarithms [514]. Long [107, 498]. Loop [418]. Loops [422]. Lorentz [446, 539]. Low [107, 18, 509, 386, 41]. Low-distortion [386]. Low-lying [509]. Lower [129, 356, 359]. Lowering [140]. Lushness [198]. Lusztig [555]. Lyapunov [346]. lying [509].

MacPherson [27]. MAD [127]. magnetic [233]. magnétiques [233]. Mahler [308, 314]. Main [200]. Manifold [549, 465]. Manifolds [210, 323, 540, 486, 243, 409, 15, 49, 54, 206, 303, 69, 14]. Manin [494]. Many [226]. Map [363, 193]. Mappings [410, 90, 145, 146]. Maps [346, 310, 534, 473, 311, 18, 179, 224, 352, 333, 286, 453, 380, 48]. Marcinkiewicz [527]. Matrices [335, 534, 517, 529, 454, 9]. Matrix [416, 454]. Matroid [61]. matter [512, 513, 521, 522, 532, 533, 541, 542, 551, 552, 563, 564]. Maximal [299, 175, 498, 485, 451, 343, 71]. Maximum [334]. meagreness [254]. Mean [130, 226]. Measure [345, 270, 268, 162, 180, 295]. Measures [412, 265, 419, 385, 482, 314, 350, 42, 71]. Metaplectic [95, 484, 153, 361]. métaplectique [149]. Method [560]. Metric [270, 93, 268, 162, 340, 386, 194]. Metrics [296, 3, 67]. Metrized [528, 475]. Meyer [213]. Mice [279]. Middle [364]. Milnor [123]. Minimal [155, 289, 129, 447, 267, 256, 331, 63]. Minkowski [439]. Minors [158]. minuscule [62]. Mirabolic [171]. Mirror [364]. Mixed [397]. Mixed-norm [397]. Mock [164]. Mod [530, 464, 491]. Model [155]. Models [302, 383, 225, 465]. Modular [554, 121, 388, 249, 96, 432, 491, 332, 152, 314, 200]. Module [349]. Modules [2, 368, 548, 207, 22, 138, 29]. Moduli [281, 373, 101, 103, 55, 57, 287, 530]. Modulo [140]. Modulus [483]. Moeglin [436]. Moments [97, 134, 58]. Monge [20, 399, 14]. Monochromatic [526]. Monodromy [395, 160]. Monoids [195, 262]. Monopoles [392]. Monotone [559, 230, 379]. Monotonically [90]. Mordell [535]. Mori [425]. Morrey [139]. Motion [158, 315]. Motives [483]. Motivic [345]. Mullen [501]. Multidimensional [276]. Multilinear [197]. Multiloop [360].

Multimarginal [311]. Multiple [83, 314, 9]. Multiplication [354, 480, 305, 24]. Multiplicative [47, 499]. Multiplier [404, 142, 415]. Multipliers [102, 365, 197, 527, 115, 294, 204]. Multiply [472]. Multivariate [495, 209]. Mumford [371]. mutation [76]. Mutually [12, 173].

Nakajima [48]. Natural [128]. nature [78]. naturels [128]. Near [185, 233]. Nearly [187, 517, 271, 272]. Negative [135]. Neumann [41, 487, 399, 10, 147]. neutral [7]. Nevanlinna [404]. Nilpotent [269]. Nilradical [444]. Nirenberg [59]. NLS [285]. Noise [170]. nombre [376]. Non [264, 186, 156, 157, 439, 348, 228, 102, 121, 254, 482, 162, 369, 406, 260, 279, 116, 360, 387, 242, 343, 89, 434, 71]. Non-Archimedean [406, 116]. Non-associative [228]. Non-Commutative [102, 264, 348]. Non-cylindrical [260]. non-doubling [71]. Non-Existence [121]. Non-homogeneous [162]. Non-local [439]. Non-meagreness [254]. Non-reductive [157]. Non-regular [482]. non-separable [89]. Non-simple [156]. Non-split [369]. Non-stable [360]. Non-tame [279]. Non-tangential [343]. Non-vanishing [186, 387]. Non-wandering [242]. Nonabelian [126]. noncommutative [29]. Noncommuting [307, 390]. Nondivergence [136]. Nonlinear [486]. Nonlinearity [434]. Nonlocal [374]. Nonnegative [301]. Nonself [218]. Nonself-adjoint [218]. Nonunital [326]. Nori [483]. Norm [397, 506]. Normal [130, 492]. normalisés [68]. Normality [175, 457]. normalized [68]. Normed [413, 207]. Norms [431]. Note [181]. Notion [405]. Nuclear [94]. Null [254]. Number [174, 372, 317, 236, 338, 249, 376, 460]. Numbers [281, 230, 476, 143]. Numerical [198].

Obstacles [177]. Obstructions [318, 293]. Obtained [181]. Octics [101]. Odd [18, 331]. ODEs [16, 1]. Odometers [510]. Oeljeklaus [409]. Omega [153]. Omega-regular [153]. Omitting [406]. One [438, 311, 92, 95, 407, 475, 98]. One-dimensional [311, 407, 98]. One-Parameter [92]. One-relation [438]. Oort [470]. Open [242, 89]. Operads [438]. opérateurs [68]. Operations [429]. Operator [397, 496, 125, 451, 14]. Operators [319, 558, 497, 516, 403, 384, 19, 104, 231, 162, 179, 56, 135, 390, 453, 444, 343, 464, 456, 68, 233, 24, 45, 71]. Optimal [311, 192, 163, 400, 169]. Optimization [374]. Orbifold [34, 433]. Orbit [175, 550]. Orbital [474, 385, 241]. orbites [300]. Orbits [535, 442, 181, 300]. Order [129, 119, 1, 205, 515, 490, 453, 212, 435, 15, 16]. Ordered [266]. Orderings [246]. Ordinary [396]. Oriented [545]. Orthogonal [335, 1, 209, 16]. Orthonormal [449]. Osculating [407]. Outer [342, 467]. outliers [454]. Overconvergent [332]. Overlaps [98].

Packets [436]. Padé [394]. pair [282]. pair-epsilon [282]. paire [282]. Pairing [488]. Pairs [428, 496, 549, 73]. Paper [202]. Parabolic [373, 539, 260, 136, 13]. paracompactness [457]. Parameter [92]. Parameters [113, 68]. paramètres [68]. Parametrization [289, 436]. Parametrized [438]. Partial [479, 529, 342, 171, 235]. partially [28]. Path [159]. Pathological [357]. Pauli [233]. PEL [396]. Pencil [356]. Percolation [504]. Perfect [246, 443]. Periodic [1, 461, 142, 16]. Periods [471]. Permutation [208]. Permutations [372]. Perturbation [260, 10]. Perturbations [107]. Pezzo [133]. Phelps [456, 558, 257]. Phenomena [357]. Picard [133]. Pick [404]. Pisot [143]. Planar [118, 347].

Plancherel [265]. Plane [289, 315]. Plans [163]. Player [127]. Poincaré [356]. Point [473, 199]. Pointed [124]. Points [309, 110, 53, 238, 93, 176, 221, 159, 180, 369, 287, 393, 128, 514]. Pointwise [473, 558, 540]. Poisson [166, 64]. Poles [547]. Polydisk [179]. Polygons [105]. Polyhedra [274]. polyhedron [44]. Polynomial [148, 556, 267, 459, 192, 353, 81]. Polynomials [337, 554, 317, 555, 1, 383, 431, 142, 223, 209, 530, 393, 16, 80]. Polytope [61]. Polytopes [120, 133, 561, 443]. pondéré [149]. Pontryagin [334]. Posedness [184]. Positive [445, 415, 391, 453, 3, 28]. Positively [217]. Post [377]. Post-critically [377]. Potential [117, 315, 4]. Potentials [100]. pour [300, 252, 149, 401]. Power [337, 434, 116]. Powers [535, 412, 140, 221, 135]. Prehomogeneous [444]. Prékopa [382]. près [233]. Prescribed [317, 166, 226]. Presented [543]. Preservers [39, 426]. Preserving [534, 286, 453]. presheaves [29]. Prime [554, 381, 140, 367]. Primes [18]. Principal [5, 364, 8, 46]. Principle [334, 15, 371]. Probabilistic [174]. Probabilities [482]. Problem [437, 301, 236, 430, 404, 273, 399, 518, 136, 226, 461]. Problems [374, 407, 260, 183, 380]. Processes [359, 76]. Product [262, 202, 312, 422, 531, 154]. Products [495, 546, 565, 218, 94, 208, 329, 510, 361]. Program [425]. Projections [156, 481]. Projective [224]. Projectivity [321, 207]. Projectors [7]. Proof [212, 501]. Proofs [202]. Propagator [285]. Properties [455, 473, 238, 352, 199, 23]. Property [495, 355, 558, 198, 329, 257, 211, 452, 300, 456, 87]. propriété [300]. Proximity [93]. pseudo [252, 43]. pseudo-amenability [43]. pseudo-coefficient [252]. Pseudocompact [528]. Pseudoconvex [250]. Pseudoharmonic [556]. Pseudolocality [74]. Pseudopowers [32]. Pseudoprime [131]. Pure [434, 154]. Purely [202].

Quadratic [557, 368, 275, 284, 23]. Quandle [132]. Quantitative [419]. Quantum [158, 62, 440, 365, 480, 231, 232, 278, 277, 22, 391, 380]. Quartic [423, 511, 567]. Quasi [543, 39, 539]. Quasi-Commutativity [39]. Quasi-homomorphisms [543]. Quasi-linear [539]. Quasianalytic [145, 463]. Quasimap [201]. Quasisplit [436]. Quasisymmetric [403]. Quaternion [388]. Quaternionic [354, 176, 491]. Queer [363]. Quintic [96]. Quite [420]. Quivers [298, 175]. Quotient [510]. Quotients [505, 367, 414, 389, 400, 500, 201].

Radius [534]. Radon [269]. Raising [152]. Rallis [547, 401]. Ramanujan [186, 121]. Ramified [5]. Randers [130, 188]. Random [504, 165, 158, 348, 253, 283, 529]. Randomly [199]. Range [107, 439, 497]. Rank [246, 349, 120, 18, 95, 475, 134, 442, 99, 530, 60, 31, 57, 9]. Ranks [189]. Rate [256, 359, 350]. Rational [362, 289, 465, 535, 190, 368, 287, 514]. Rationality [172]. rationnels [514]. Real [215, 281, 101, 489, 112, 442, 31]. Rearrangement [439, 198]. recombination [76]. Rectifiability [163]. Recurrence [192, 83]. Reduced [546, 565]. Reducibility [95, 35, 46]. reducing [24]. Reduction [408, 400]. Reductions [349, 131, 470]. Reductive [288, 368, 157]. Rees [289]. Refinement [382]. Reflection [298, 254]. reflexive [57]. Reflexivity [495]. Regular [438, 525, 221, 482, 153]. Regularity [214]. Regularization [437, 216, 273]. Regularized [184]. régulateurs [376]. Regulators [471, 423, 376]. Related [185, 383, 374, 431, 168, 380, 243, 20, 59]. relation [438]. Relations [292]. Relative [500, 469]. Remark [261]. renormings [33]. Repeated [411]. Representation [337, 506, 361, 477]. Representations [155, 288, 5, 265, 474, 458, 84, 113, 408, 287, 500, 531, 86, 562]. Representing [102]. Repulsive [311]. Residues [444]. Resolution [447]. Resonance [110]. Resonances [233, 233]. Resonant [117]. Respect [407]. restrictions [86]. Results [184, 243, 20, 21]. Reversibility [76]. Revisiting [48]. Revolution [297]. Rham [477]. Ricci [3, 157, 51, 74, 409]. Riemann [51, 109, 330]. Riemannian [15, 217]. Riesz [129, 439]. Right [372]. Rigidity [377, 274, 20]. Ring [17]. Rings [342]. Robinson [171]. Roch [330]. Root [476]. roots [255]. Rotation [130, 328, 313]. rotund [33]. Rough [384, 183]. Roughening [169]. row [398]. Rudin [492]. Ruelle [550]. ruled [60]. Rules [5].

S. [52]. Salem [412, 143]. SANE [127]. Saturated [475, 127]. scalar [67]. Scheme [227]. Schemes [111]. Schensted [171]. Schiffmann [547]. Schoenberg [329]. Schrödinger [285, 540]. Schubert [407]. Schur [264, 115, 530, 261]. Schwartz [489]. Second [1, 402, 274, 15, 16]. Segal [43]. Seifert [248]. Selberg [26]. selection [76]. Self [18, 69]. self-equivalences [69]. Self-Maps [18]. Selmer [275, 322, 152]. Semi [161, 62]. Semi-Dirichlet [161]. Semi-simplicity [62]. semicircular [223]. Semicrossed [495, 218]. Semigroup [427, 455]. Semigroups [218]. semilinear [4]. Semisimple [288, 31]. separable [89]. Sequence [207, 492]. Sequences [449, 525, 492, 451, 83]. série [252]. Series [5, 109, 559, 528, 475, 116, 500, 252, 86, 8, 46]. Set [215, 213]. Sets [337, 238, 254, 498, 199, 194, 242, 77]. seuils [233]. seven [443]. Several [307]. Shafarevich [432]. Shapiro [492]. Sharp [397, 294]. Sheaf [549]. sheaves [29, 57].

Shifted [499]. Shilov [64]. Shimura [396, 103, 176, 292, 287, 280, 234, 36, 470]. Short [165]. Shtukas [548]. Shuffle [159]. Siegel [332]. Sierpinski [79, 30]. sieve [30]. Signed [275]. Signed-Selmer [275]. Similar [118, 492]. Similarity [302, 19]. Simple [202, 546, 565, 150, 562, 151, 442, 156]. Simplest [423]. Simplices [443]. Simplicial [89]. simplicity [62]. Simultaneous [411]. Singular [139, 103, 384, 539, 392, 45, 461]. Skew [335, 342, 530]. Skew-Hamiltonian [335]. Skolem [308]. Smale [550]. Small [543, 120, 55]. Smooth [129, 55, 459, 15]. Smoothing [528, 475]. Smoothness [197]. Sobolev [4, 237, 351, 435]. Soergel [562]. Sofic [312]. Solitar [137]. Solitons [157, 117]. Solutions [4, 347, 526, 459, 539, 183, 434, 226, 540, 461]. Solvability [260, 136]. Solvable [524]. Some [148, 524, 374, 20, 10, 434, 67, 70]. Sommes [320]. Souslin [519]. Space [310, 130, 301, 101, 430, 105, 224, 561, 286, 530, 241, 128, 256, 467, 24]. Spaces [446, 523, 281, 478, 373, 157, 397, 139, 472, 496, 237, 102, 413, 217, 188, 270, 253, 93, 458, 497, 268, 527, 404, 104, 271, 162, 441, 198, 330, 168, 537, 331, 550, 144, 386, 194, 116, 531, 295, 343, 62, 6, 457, 7, 57, 33, 89, 42, 469, 25, 204, 272]. Spanning [553]. Spatial [132]. Special [1, 11, 234, 503, 16, 78]. Species [195]. Specified [142]. Specifying [159]. Spectra [304, 213]. Spectral [534, 326, 516, 329, 98]. Spectrum [17, 453]. Sphere [224, 461, 273, 45]. spheres [248]. Spherical [155, 484, 306]. Splayed [344]. Split [108, 368, 112, 369]. Splitting [420, 112, 150]. Springer [398]. Square [337, 502]. Square-free [502]. Squared [217]. Squares [492, 507]. Stability [410, 337, 346, 347, 439, 469]. Stable [530, 360]. Stably [19]. stacks [345]. Stahl [52]. Standard [547]. Standing [214].

Stationary [170]. Steiner [21]. Steklov [518]. Step [269]. Stochastic [170]. strata [470]. Strict [415]. Strictly [250]. String [537]. Strong [236, 351, 77, 394]. Strongly [355, 375, 19]. Structure [523, 403, 234, 364]. Structures [245, 106, 340]. Sturmian [300]. sturmiennes [300]. Subdivisions [61]. Subelliptic [183]. Subgroup [479]. Subgroups [499]. Subharmonic [351]. Sublattices [118]. Submanifolds [259]. subpolyhedron [465]. Subsequences [174]. Subsequent [202]. subsets [89]. Subsolutions [216]. Subspaces [286, 116, 24]. Sum [9, 83]. Summable [355, 516]. Sums [355, 55, 499]. Sumset [327]. Sumsets [424, 115]. Superalgebras [363]. Supercategory [545]. Supercuspidal [265, 408]. Superharmonic [478]. supersingular [36]. Support [238, 482]. Surface [91, 109, 414, 481, 530, 511, 128, 63, 567]. Surfaces [437, 395, 354, 318, 130, 448, 414, 356, 352, 133, 210, 485, 193, 225, 520, 60, 49]. Symbolic [221]. Symbols [390, 81]. Symmetric [264, 217, 490, 525, 425, 315, 359, 306, 6]. Symmetries [100, 429]. Symmetry [364]. Symplectic [245, 290, 201]. Synthesis [329]. System [184]. Systems [373, 347, 496, 339, 177, 191, 331, 98, 341, 235, 394, 4, 21].

tame [279]. Tameness [215]. Tandori [523]. tangential [343]. Tannakian [427]. Tate [432, 305]. Taut [248]. Tensor [284, 361, 51]. Terms [185, 129, 486]. Ternary [459]. test [247]. Their [508, 506, 402, 21, 406, 469]. Theorem [195, 174, 325, 308, 545, 472, 122, 338, 167, 421, 494, 243, 126]. Theorems [373, 473, 82]. theoretic [352]. Theory [174, 263, 216, 318, 1, 416, 144, 538, 200, 28, 16, 266, 240]. Theta [95, 164, 408, 464]. Three [273, 168, 210, 303]. Three-manifolds [303]. Three-sphere [273]. Threefold [96]. Threefolds [65, 225]. thresholds [233]. Tietze [48]. Tight [182]. Tilings [413]. time [339]. Titchmarsh [560]. Toeplitz [497, 390]. Toma [409]. Topological [94, 144, 452, 173]. Topology [160, 537, 538]. Torelli [373]. Toric [383, 105, 65, 467, 323, 400]. torsion [51, 222]. torsion-free [51]. Torsors [124]. Torus [429, 22, 293]. Touch [159]. towers [460, 53]. Trace [231, 313, 151]. traces [401]. Trajectories [220]. Transcendental [78]. Transfer [265, 474, 115, 252]. Transference [204]. Transfert [252]. Transform [269, 64]. Transformation [240, 316]. Transitive [562]. Transitivity [517]. Translation [34, 433]. Transport [311, 419]. Transportation [163, 350]. Transposes [529]. treatment [40]. Triangles [137]. Triangulated [105]. Triangulations [504]. Trichotomy [339]. Trigonometric [559]. Trinomials [499]. Triple [219, 21]. Triples [326]. Trivial [355]. Tropical [362, 318, 190, 160, 323]. Tunneling [117]. Tunnels [395]. Tutte [223]. Twin [554]. Twisted [319, 140, 284, 418]. Twists [280, 23]. Two [108, 348, 269, 536, 562, 293, 152, 530, 212, 500, 225, 60, 31]. Two-color [562]. Two-dimensional [536]. Two-faced [348]. Type [396, 187, 398, 374, 448, 482, 528, 356, 399, 56, 333, 270, 253, 276, 469, 470]. Types [291, 291].

Ultrafilters [355]. ultraspherical [80]. Umbilical [259]. Unanimity [185]. Uniform [553, 165, 559, 257]. Uniformity [508]. Uniformly [229, 33, 42]. Union [263, 355]. Unipotent [408]. Unique [37, 151, 279]. Uniquely [178]. Uniqueness [487]. Unit [317]. unitaires [401]. Unital [94, 566, 468]. Unitary [309, 130, 319, 265, 302, 442, 153, 234]. Universal [109, 190, 243]. Universality [209]. Unknotting [395]. Unperforated [496]. Urysohn [194].

Valuations [61, 393]. Value [1, 16]. valued [100, 390, 204]. Values [284, 502, 78, 11, 42, 23, 314, 387, 58]. vanishing [186, 387]. Variable [446]. Variables [307, 348]. variant [85]. variante [401]. variational [15]. Variations [239, 364, 63, 30]. variétés [514]. Varieties [299, 185, 396, 489, 290, 176, 267, 249, 160, 168, 292, 305, 234, 477, 322, 467, 27, 470]. Variety [536, 36]. Various [352]. Varying [201]. Vector [281, 534, 130, 168, 60, 204]. vector-valued [204]. vectorial [83]. Vectors [119, 555]. Verma [22]. Version [207]. Versus [221, 457]. Vertex [319]. very [60]. Via [318, 95, 143, 24, 71]. Vilkovisky [537]. Vinogradov [276]. Vinogradov-type [276]. Viot [76]. Vision [227]. Vojta [75]. Volume [148, 243, 512, 513, 521, 522, 532, 533, 541, 542, 551, 552, 563, 564].

Walks [165, 158, 283]. wandering [242]. Wasserstein [441]. Wavelets [99]. Waves [214]. Weak [216, 302, 515, 211, 452, 539, 183, 371]. Weakly [229, 177, 182, 6]. Wedge [429]. Weight [152]. Weighted [397, 509, 295, 204]. Weights [478, 464, 358, 25]. Weil [122]. Weingarten [448]. Well [184]. Well-Posedness [184]. Weyl [428, 301]. Whitney [270]. Whitney-type [270]. Whittaker [31, 306]. Wishart [529]. without [482]. Witten [323]. Wronskian [212].

Yau [505, 556, 96].

Zero [534, 230, 442, 452, 47]. zero-divisors [47]. Zeros [317, 509]. zeta [83]. zetafunctions [83]. Zygmund [162].

References

Chiang:2006:VDT

[1] Yik-Man Chiang and Mourad E. H. Ismail. On value distribution theory of second order periodic ODEs, special functions and orthogonal polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 58 (4):726–767, August 2006. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [16].

Bell:2009:MAI

[2] J. P. Bell and K. G. Hare. On Z-modules of algebraic integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 61(??):264–281, ???? 2009. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [138].

Anchouche:2010:ABC

[3] Boudjemâa Anchouche. On the asymptotic behavior of complete Kähler metrics of positive Ricci curvature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1): 3–18, February 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bouchekif:2010:SSE

[4] Mohammed Bouchekif and Yasmina Nasri. Solutions for semilinear elliptic systems with critical Sobolev exponent and Hardy potential. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1):19–33, February 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Campbell:2010:BRR

[5] Peter S. Campbell and Monica Nevins. Branching rules for ramified principal series representations of GL(3) over a p-adic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1):34–51, February 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deng:2010:AAW

[6] Shaoqiang Deng. An algebraic approach to weakly symmetric Finsler spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (1):52–73, February 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ducrot:2010:PGE

[7] Arnaud Ducrot, Zhihua Liu, and Pierre Magal. Projectors on the generalized eigenspaces for neutral functional differential equations in L^{p} spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1): 74–93, February 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kuo:2010:LCG

[8] Wentang Kuo. The Langlands correspondence on the generic irreducible constituents of principal series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1): 94–108, February 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Li:2010:SHM

[9] Chi-Kwong Li and Yiu-Tung Poon. Sum of Hermitian matrices with given eigenvalues: inertia, rank, and multiple eigenvalues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1):109–132, February 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Makarov:2010:SAP

[10] Konstantin A. Makarov and Anna Skripka. Some applications of the perturbation determinant in finite von Neumann algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1):133–156, February 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Masri:2010:SVC

[11] Riad Masri. Special values of class group L-functions for CM fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1): 157–181, February 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Prajs:2010:MAD

[12] Janusz R. Prajs. Mutually aposyndetic decomposition of homogeneous continua. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1):182–201, February 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tang:2010:IEP

[13] Lin Tang. Interior h^{1} estimates for parabolic equations with LMO coefficients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1):202–217, February 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Xing:2010:GDC

[14] Yang Xing. The general definition of the complex Monge–Ampère operator on compact Kähler manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(1): 218–239, February 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Azagra:2010:SOS

[15] Daniel Azagra and Robb Fry. A second order smooth variational principle on Riemannian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):241–260, April 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chiang:2010:EVD

[16] Yik-Man Chiang and Mourad E. H. Ismail. Erratum to: On value distribution theory of second order periodic ODEs, special functions and orthogonal polynomials [Cno. 2245272]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):261, April 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [1].

Goresky:2010:SEC

[17] Mark Goresky and Robert MacPherson. On the spectrum of the equivariant cohomology ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):262–283, April 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grbic:2010:SML

[18] Jelena Grbić and Stephen Theriault. Self-maps of low rank Lie groups at odd primes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):284–304, April 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

He:2010:ASC

[19] Hua He, Yunbai Dong, and Xianzhou Guo. Approximation and similarity classification of stably finitely strongly irreducible decomposable operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2): 305–319, April 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jerrard:2010:SRR

[20] Robert L. Jerrard. Some rigidity results related to Monge–Ampère functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):320–354, April 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kral:2010:CRS

[21] Daniel Král, Edita Máčajová, Attila Pór, and Jean-Sébastien Sereni. Characterisation results for Steiner triple systems and their application to edgecolourings of cubic graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2): 355–381, April 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lu:2010:VMQ

[22] Rencai Lü and Kaiming Zhao. Verma modules over quantum torus Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):382–399, April 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Prasanna:2010:APC

[23] Kartik Prasanna. On p-adic properties of central L-values of quadratic twists of an elliptic curve. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):400–414, April 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sun:2010:CRS

[24] Shunhua Sun, Dechao Zheng, and Changyong Zhong. Classification of reducing subspaces of a class of multiplication operators on the Bergman space via the Hardy space of the bidisk. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2): 415–438, April 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sundhall:2010:HFH

[25] Marcus Sundhäll and Edgar Tchoundja. On Hankel forms of higher weights: the case of Hardy spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2):439–455, April 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yang:2010:CSF

[26] Tonghai Yang. The Chowla–Selberg formula and the Colmez conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2): 456–472, April 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yun:2010:GMC

[27] Zhiwei Yun. Goresky–MacPherson calculus for the affine flag varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(2): 473–480, April 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Casals-Ruiz:2010:EAG

[28] Montserrat Casals-Ruiz and Ilya V. Kazachkov. Elements of algebraic geometry and the positive theory of partially commutative groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(3):481–519, June 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Eriksen:2010:CND

[29] Eivind Eriksen. Computing noncommutative deformations of presheaves and sheaves of modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(3):520–542, June 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hare:2010:MVS

[30] Kevin G. Hare. More variations on the Sierpiński sieve. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(3):543–562, June 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ishii:2010:WFR

[31] Taku Ishii. Whittaker functions on real semisimple Lie groups of rank two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (3):563–581, June 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Konyagin:2010:DP

[32] Sergei V. Konyagin, Carl Pomerance, and Igor E. Shparlinski. On the distribution of pseudopowers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(3):582–594, June 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martinez:2010:LUR

[33] J. F. Martínez, A. Moltó, J. Orihuela, and S. Troyanski. On locally uniformly rotund renormings in C(K) spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (3):595–613, June 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pronk:2010:TGO

[34] Dorette Pronk and Laura Scull. Translation groupoids and orbifold cohomology. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(3):614–645, June 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See erratum [433].

Rupp:2010:R

[35] R. Rupp and A. Sasane. Reducibility in A_{R}(K), C_{R}(K), and A(K). Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(3):

646–667, June 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vollaard:2010:SLS

[36] Inken Vollaard. The supersingular locus of the Shimura variety for GU(1, s). Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (3):668–720, June 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boocher:2010:FFU

[37] Adam Boocher, Michael Daub, Ryan K. Johnson, H. Lindo, S. Loepp, and Paul A. Woodard. Formal fibers of unique factorization domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4): 721–736, August 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ditzian:2010:ADA

[38] Z. Ditzian and A. Prymak. Approximation by dilated averages and K-functionals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4):737–757, August 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dolinar:2010:GPQ

[39] Gregor Dolinar and Bojan Kuzma. General preservers of quasi-commutativity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (4):758–786, August 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Landquist:2010:ETC

[40] E. Landquist, P. Rozenhart, R. Scheidler, J. Webster, and Q. Wu. An explicit treatment of cubic function fields with applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4):787–807, August 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Legendre:2010:ELE

[41] Eveline Legendre. Extrema of low eigenvalues of the Dirichlet–Neumann Laplacian on a disk. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4):808–826, August 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ouyang:2010:BFC

[42] Caiheng Ouyang and Quanhua Xu. BMO functions and Carleson measures with values in uniformly convex spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (4):827–844, August 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Samei:2010:BPA

[43] Ebrahim Samei, Nico Spronk, and Ross Stokke. Biflatness and pseudoamenability of Segal algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4): 845–869, August 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Valdimarsson:2010:BLP

[44] Stefán Ingi Valdimarsson. The Brascamp–Lieb polyhedron. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4): 870–888, August 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Xia:2010:SIO

[45] Jingbo Xia. Singular integral operators and essential commutativity on the sphere. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4):889–913, August 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zorn:2010:RPS

[46] Christian Zorn. Reducibility of the principal series for S̃p_{2}(F) over a p-adic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(4):914–960, August 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aleman:2010:MII

[47] Alexandru Aleman, Peter Duren, María J. Martín, and Dragan Vukotić. Multiplicative isometries and isometric zero-divisors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5):961–974, October 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bjorndahl:2010:RTN

[48] Christina Bjorndahl and Yael Karshon. Revisiting Tietze–Nakajima: local and global convexity for maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5): 975–993, October 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Breslin:2010:CBS

[49] William Breslin. Curvature bounds for surfaces in hyperbolic 3-manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (5):994–1010, October 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Buckingham:2010:FCF

[50] Paul Buckingham and Victor Snaith. Functoriality of the canonical fractional Galois ideal. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5):1011–1036, October 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Calvino-Louzao:2010:RET

[51] E. Calviño-Louzao, E. García-Río, and R. Vázquez-Lorenzo. Riemann extensions of torsion-free connections with degenerate Ricci tensor. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5):1037– 1057, October 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2010:CS

[52] Yichao Chen and Yanpei Liu. On a conjecture of S. Stahl. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5): 1058–1059, October 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Darmon:2010:HPT

[53] Henri Darmon and Ye Tian. Heegner points over Towers of Kummer extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5):1060–1081, October 2010. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Godinho:2010:FGM

[54] Leonor Godinho and M. E. Sousa-Dias. The fundamental group of S^{1}-manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (5):1082–1098, October 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldmakher:2010:CSS

[55] Leo Goldmakher. Character sums to smooth moduli are small. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5): 1099–1115, October 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jin:2010:DLO

[56] Yongyang Jin and Genkai Zhang. Degenerate p-Laplacian operators and Hardy type inequalities on H-type groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5):1116–1130, October 2010. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kleppe:2010:MSR

[57] Jan O. Kleppe. Moduli spaces of reflexive sheaves of rank 2. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5): 1131–1154, October 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Young:2010:MCV

[58] Matthew P. Young. Moments of the critical values of families of elliptic curves, with applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5): 1155–1181, October 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yue:2010:FFR

[59] Hong Yue. A fractal function related to the John–Nirenberg inequality for Q_{α}(R^{n}). Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(5):1182–1200, October 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alzati:2010:CVA

[60] Alberto Alzati and Gian Mario Besana. Criteria for very ampleness of rank two vector bundles over ruled surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (6):1201–1227, December 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ardila:2010:VMP

[61] Federico Ardila, Alex Fink, and Felipe Rincón. Valuations for matroid polytope subdivisions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1228– 1245, December 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chaput:2010:QCM

[62] P. E. Chaput, L. Manivel, and N. Perrin. Quantum cohomology of minuscule homogeneous spaces III. Semisimplicity and consequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1246– 1263, December 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2010:HVM

[63] Jingyi Chen and Ailana Fraser. Holomorphic variations of minimal disks with boundary on a Lagrangian surface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (6):1264–1275, December 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

ElWassouli:2010:GPT

[64] Fouzia El Wassouli. A generalized Poisson transform of an L^{p}-function over the Shilov boundary of the ndimensional Lie ball. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1276– 1292, December 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kasprzyk:2010:CTF

[65] Alexander M. Kasprzyk. Canonical toric Fano threefolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1293– 1309, December 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:2010:IHA

[66] Kyu-Hwan Lee. Iwahori–Hecke algebras of SL_{2} over 2-dimensional local fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62 (6):1310–1324, December 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mo:2010:SEC

[67] Xiaohuan Mo and Changtao Yu. On some explicit constructions of Finsler metrics with scalar flag curvature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1325– 1339, December 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moeglin:2010:HOE

[68] C. Mœglin. Holomorphie des opérateurs d’entrelacement normalisés à l’aide des paramètres d’Arthur. (French) [Holomorphism of normalized interlacing operators with the help of Arthur parameters]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1340–1386, December 2010. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pamuk:2010:HSE

[69] Mehmetcik Pamuk. Homotopy selfequivalences of 4-manifolds with free fundamental group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1387– 1403, December 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Saroglou:2010:CES

[70] Christos Saroglou. Characterizations of extremals for some functionals on convex bodies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1404–1418, December 2010. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yang:2010:BEM

[71] Dachun Yang and Dongyong Yang. BMO-estimates for maximal operators via approximations of the identity with non-doubling measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 62(6):1419– 1434, December 2010. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Banica:2011:FBL

[72] T. Banica, S. T. Belinschi, M. Capitaine, and B. Collins. Free Bessel laws. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1): 3–37, February 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brudern:2011:AFP

[73] Jörg Brüdern and Trevor D. Wooley. Asymptotic formulae for pairs of diagonal cubic equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1):38–54, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chau:2011:PRF

[74] Albert Chau, Luen-Fai Tam, and Chengjie Yu. Pseudolocality for the Ricci flow and applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1): 55–85, February 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2011:VC

[75] Xi Chen. On Vojta’s 1 + ϵ conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1):86–103, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feng:2011:RIF

[76] Shui Feng, Byron Schmuland, Jean Vaillancourt, and Xiaowen Zhou. Reversibility of interacting Fleming–Viot processes with mutation, selection, and recombination. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1):104–122, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Granirer:2011:SES

[77] Edmond E. Granirer. Strong and extremely strong Ditkin sets for the Banach algebras A_{p}^{r}(G) = A_{p}∩L^{r}(G). Canadian Journal of Mathematics = Journal canadien de mathématiques, 63 (1):123–135, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gun:2011:TNS

[78] Sanoli Gun, M. Ram Murty, and Purusottam Rath. Transcendental nature of special values of L-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1): 136–152, February 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hambly:2011:AFA

[79] B. M. Hambly. Asymptotics for functions associated with heat flow on the Sierpinski carpet. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1):153–180, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ismail:2011:CCD

[80] Mourad E. H. Ismail and Josef Obermaier. Characterizations of continuous and discrete q-ultraspherical polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1):181–199, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rahman:2011:EPE

[81] Mizan Rahman. An explicit polynomial expression for a q-analogue of the 9-j symbols. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(1):200–221, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wang:2011:LTA

[82] Jiun-Chau Wang. Limit theorems for additive conditionally free convolution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63 (1):222–240, February 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Essouabri:2011:MZF

[83] Driss Essouabri, Kohji Matsumoto, and Hirofumi Tsumura. Multiple zetafunctions associated with linear recurrence sequences and the vectorial sum formula. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):241–276, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ghate:2011:LIG

[84] Eknath Ghate and Vinayak Vatsal. Locally indecomposable Galois representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):277–297, April 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gun:2011:VLC

[85] Sanoli Gun and V. Kumar Murty. A variant of Lehmer’s conjecture, II: the CM-case. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):298–326, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jantzen:2011:DSA

[86] Chris Jantzen. Discrete series for padic SO(2n) and restrictions of representations of O(2n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):327–380, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ji:2011:CCA

[87] Kui Ji and Chunlan Jiang. A complete classification of AI algebras with the ideal property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):381–412, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Konvalinka:2011:GFH

[88] Matjaž Konvalinka and Mark Skandera. Generating functions for Hecke algebra characters. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):413–435, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mine:2011:SCO

[89] Kotaro Mine and Katsuro Sakai. Simplicial complexes and open subsets of nonseparable LF-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):436–459, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pavlicek:2011:MCM

[90] Libor Pavlíček. Monotonically controlled mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(2):460–480, April 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baragar:2011:ACK

[91] Arthur Baragar. The ample cone for a K3 surface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(3):481–499, June 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dadarlat:2011:OPC

[92] Marius Dadarlat, George A. Elliott, and Zhuang Niu. One-parameter continuous fields of Kirchberg algebras. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63 (3):500–532, June 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Espinola:2011:BPP

[93] Rafa Espínola and Aurora Fernández-León. On best proximity points in metric and Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(3):533–550, June 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hadwin:2011:TFE

[94] Don Hadwin, Qihui Li, and Junhao Shen. Topological free entropy dimensions in nuclear C^{∗}-algebras and in full free products of unital C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63 (3):551–590, June 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hanzer:2011:ROR

[95] Marcela Hanzer and Goran Muić. Rank one reducibility for metaplectic groups via theta correspondence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(3):591–615, June 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:2011:MQC

[96] Edward Lee. A modular quintic Calabi– Yau threefold of level 55. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(3): 616–633, June 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lu:2011:HMF

[97] Guangshi Lü. On higher moments of Fourier coefficients of holomorphic cusp forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(3):634–647, June 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ngai:2011:SAL

[98] Sze-Man Ngai. Spectral asymptotics of Laplacians associated with onedimensional iterated function systems with overlaps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(3):648–688, June 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Olphert:2011:HRW

[99] Sean Olphert and Stephen C. Power. Higher rank wavelets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(3):689–720, June 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Autin:2011:ICV

[100] Aymeric Autin. Isoresonant complexvalued potentials and symmetries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63 (4):721–754, August 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chu:2011:GMS

[101] Kenneth C. K. Chu. On the geometry of the moduli space of real binary octics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(4):755–797, August 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Daws:2011:RMF

[102] Matthew Daws. Representing multipliers of the Fourier algebra on noncommutative L^{p} spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(4):798–825, August 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Errthum:2011:SMS

[103] Eric Errthum. Singular moduli of Shimura curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(4):826–861, August 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hosokawa:2011:LCC

[104] Takuya Hosokawa, Pekka J. Nieminen, and Shûichi Ohno. Linear combinations of composition operators on the Bloch spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(4):862–877, August 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Howard:2011:TGT

[105] Benjamin Howard, Christopher Manon, and John Millson. The toric geometry of triangulated polygons in Euclidean space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(4):878–937, August 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Li-Bland:2011:ACA

[106] David Li-Bland. AV–Courant algebroids and generalized CR structures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(4):

938–960, August 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bouclet:2011:LFE

[107] Jean-Marc Bouclet. Low frequency estimates for long range perturbations in divergence form. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(5):961–991, October 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruin:2011:AGT

[108] Nils Bruin and Kevin Doerksen. The arithmetic of genus two curves with (4, 4)-split Jacobians. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(5):992–1024, October 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clouatre:2011:USR

[109] Raphaël Clouâtre. Universal series on a Riemann surface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(5):1025– 1037, October 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohen:2011:CPR

[110] D. Cohen, G. Denham, M. Falk, and A. Varchenko. Critical points and resonance of hyperplane arrangements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63 (5):1038–1057, October 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Easton:2011:CS

[111] Robert W. Easton. S_{3}-covers of schemes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63 (5):1058–1082, October 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaletha:2011:DSI

[112] Tasho Kaletha. Decomposition of splitting invariants in split real groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63 (5):1083–1106, October 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Liu:2011:GRP

[113] Baiying Liu. Genericity of representations of p-adic Sp_{2n} and local Langlands parameters. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(5):1107–1136, October 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moy:2011:DAP

[114] Allen Moy. Distribution algebras on padic groups and Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(5): 1137–1160, October 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Neuwirth:2011:TFM

[115] Stefan Neuwirth and Éric Ricard. Transfer of Fourier multipliers into Schur multipliers and sumsets in a discrete group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63 (5):1161–1187, October 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sliwa:2011:CSN

[116] Wiesław Śliwa and Agnieszka Ziemkowska.

On complemented subspaces of non-Archimedean power series spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(5): 1188–1200, October 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Salem:2011:RTF

[117] Walid K. Abou Salem and Catherine Sulem. Resonant tunneling of fast solitons through large potential barriers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63 (6):1201–1219, December 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baake:2011:SSP

[118] Michael Baake, Rudolf Scharlau, and Peter Zeiner. Similar sublattices of planar lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(6):1220–1237, December 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bump:2011:CBI

[119] Daniel Bump and Maki Nakasuji. Casselman’s basis of Iwahori vectors and the Bruhat order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(6):1238– 1253, December 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DAzevedo:2011:CCP

[120] Antonio Breda D’Azevedo, Gareth A. Jones, and Egon Schulte. Constructions of chiral polytopes of small rank. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(6):1254– 1283, December 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dewar:2011:NER

[121] Michael Dewar. Non-existence of Ramanujan congruences in modular forms of level four. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(6):1284–1306, December 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dimitrov:2011:BBW

[122] Ivan Dimitrov and Ivan Penkov. A Bott–Borel–Weil theorem for diagonal ind-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(6):1307–1327, December 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gun:2011:CCM

[123] Sanoli Gun, M. Ram Murty, and Purusottam Rath. On a conjecture of Chowla and Milnor. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(6):1328– 1344, December 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jardine:2011:PT

[124] J. F. Jardine. Pointed torsors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(6):1345– 1363, December 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Meinrenken:2011:CDO

[125] Eckhard Meinrenken. The cubic Dirac operator for infinite-dimensonal Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(6):1364–1387, December 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Misamore:2011:NEV

[126] Michael D. Misamore. Nonabelian H^{1} and the Étale Van Kampen Theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63 (6):1388–1415, December 2011. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shelah:2011:MSF

[127] Saharon Shelah. MAD saturated families and SANE player. Canadian Journal of Mathematics = Journal canadien de mathématiques, 63(6):1416– ??, December 2011. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boissiere:2012:ANE

[128] Samuel Boissière. Automorphismes naturels de l’espace de Douady de points sur une surface. (French). [Natural isomorphisms on the points in a surface in Douady space]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(1):3–23, February 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borodachov:2012:LOT

[129] S. V. Borodachov. Lower order terms of the discrete minimal Riesz energy on smooth closed curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(1):24–43, February 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carvalho:2012:SRC

[130] T. M. M. Carvalho, H. N. Moreira, and K. Tenenblat. Surfaces of rotation with constant mean curvature in the direction of a unitary normal vector field in a Randers space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(1):44–80, February 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

David:2012:PRE

[131] C. David and J. Wu. Pseudoprime reductions of elliptic curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(1): 81–101, February 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ishii:2012:QCI

[132] Atsushi Ishii and Masahide Iwakiri. Quandle cocycle invariants for spatial graphs and knotted handlebodies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(1): 102–122, February 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:2012:GPP

[133] Jae-Hyouk Lee. Gosset polytopes in Picard groups of del Pezzo surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(1): 123–150, February 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miller:2012:MRE

[134] Steven J. Miller and Siman Wong. Moments of the rank of elliptic curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64 (1):151–182, February 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nowak:2012:NPL

[135] Adam Nowak and Krzysztof Stempak.

Negative powers of Laguerre operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64 (1):183–216, February 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tang:2012:SCD

[136] Lin Tang. W_{ω}^{2}, p-solvability of the Cauchy–Dirichlet problem for nondivergence parabolic equations with BMO coefficients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(1):217–??, February 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Allcock:2012:TBS

[137] Daniel Allcock. Triangles of Baumslag– Solitar groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(2):241–??, April 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bell:2012:CMA

[138] Jason P. Bell and Kevin G. Hare. Corrigendum to “On Z-modules of Algebraic Integers”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(2):254–??, April 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [2].

Chen:2012:CCS

[139] Yanping Chen, Yong Ding, and Xinxia Wang. Compactness of commutators for singular integrals on Morrey spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64 (2):257–??, April 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dahmen:2012:LLM

[140] Sander R. Dahmen and Soroosh Yazdani. Level lowering modulo prime powers and twisted Fermat equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64 (2):282–??, April 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hurlburt:2012:HCF

[141] Chris Hurlburt and Jeffrey Lin Thunder. Hermite’s constant for function fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(2):301–??, April 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ingram:2012:CPP

[142] Patrick Ingram. Cubic polynomials with periodic cycles of a specified multiplier.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 64 (2):318–??, April 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McKee:2012:SNP

[143] James McKee and Chris Smyth. Salem numbers and Pisot numbers via interlacing. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(2):345–??, April 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Meyer:2012:ATS

[144] Ralf Meyer and Ryszard Nest. C^{∗}- algebras over topological spaces: Filtrated K-theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(2):368–??, April 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rainer:2012:LQM

[145] Armin Rainer. Lifting quasianalytic mappings over invariants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(2):409–??, April 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shafikov:2012:HMB

[146] Rasul Shafikov and Kaushal Verma.

Holomorphic mappings between domains in C^{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(2):429–??, April 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sherman:2012:CIG

[147] David Sherman. On cardinal invariants and generators for von Neumann algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(2):455–??, April 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chamorro:2012:SFI

[148] Diego Chamorro. Some functional inequalities on polynomial volume growth Lie groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(3):481–??, June 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Li:2012:LFP

[149] Wen-Wei Li. Le lemme fondamental pondéré pour le groupe métaplectique. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64 (3):497–??, June 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Li:2012:SIL

[150] Zhiqiang Li. On the simple inductive limits of splitting interval algebras with dimension drops. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(3):544–??, June 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nawata:2012:FGS

[151] Norio Nawata. Fundamental group of simple C^{∗}-algebras with unique trace III. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(3):573–??, June 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nekovar:2012:LRA

[152] Jan Nekovár. Level raising and anticyclotomic Selmer groups for Hilbert modular forms of weight two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(3): 588–??, June 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pantano:2012:GOR

[153] Alessandra Pantano, Annegret Paul, and Susana A. Salamanca-Riba. The genuine omega-regular unitary dual of the metaplectic group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(3):669–??, June 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thomsen:2012:PIC

[154] Klaus Thomsen. Pure infiniteness of the crossed product of an AH-algebra by an endomorphism. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(3):705–??, June 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Achab:2012:ABK

[155] Dehbia Achab and Jacques Faraut.

Analysis of the Brylinski–Kostant model for spherical minimal representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64 (4):721–??, August 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:2012:HCP

[156] Lawrence G. Brown and Hyun Ho Lee. Homotopy classification of projections in the Corona algebra of a non-simple C^{∗}-algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(4):755–??, August 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Calvaruso:2012:RSG

[157] Giovanni Calvaruso and Anna Fino. Ricci solitons and geometry of fourdimensional non-reductive homogeneous spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(4):778–??, August 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chapon:2012:QRW

[158] François Chapon and Manon Defosseux. Quantum random walks and minors of Hermitian Brownian motion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(4): 805–??, August 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haglund:2012:CSC

[159] J. Haglund, J. Morse, and M. Zabrocki. A compositional shuffle conjecture specifying touch points of the Dyck path. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64 (4):822–??, August 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Helm:2012:MFT

[160] David Helm and Eric Katz. Monodromy filtrations and the topology of tropical varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(4):845–??, August 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hu:2012:BSD

[161] Ze-Chun Hu and Wei Sun. Balayage of semi-Dirichlet forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(4):869–??, August 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hytonen:2012:BCZ

[162] Tuomas Hytönen, Suile Liu, Dachun Yang, and Dongong Yang. Boundedness of Calderón–Zygmund operators on nonhomogeneous metric measure spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64 (4):892–??, August 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCann:2012:ROT

[163] Robert J. McCann, Brendan Pass, and Micah Warren. Rectifiability of optimal transportation plans. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(4):924–??, August 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McIntosh:2012:HKF

[164] Richard J. McIntosh. The H and K families of mock theta functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(4): 935–??, August 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:2012:DSU

[165] Jonathan M. Borwein, Armin Straub, James Wan, and Wadim Zudilin. Densities of short uniform random walks. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64 (5):961–??, October 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Damianou:2012:PBP

[166] Pantelis A. Damianou and Fani Petalidou. Poisson brackets with prescribed Casimirs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(5):991–??, October 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fiorilli:2012:TBF

[167] Daniel Fiorilli. On a theorem of Bombieri, Friedlander, and Iwaniec. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64 (5):1019–??, October 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koh:2012:HAR

[168] Doowon Koh and Chun-Yen Shen. Harmonic analysis related to homogeneous varieties in three dimensional vector spaces over finite fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(5): 1036–??, October 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Plakhov:2012:ORC

[169] Alexander Plakhov. Optimal roughening of convex bodies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(5):1058–??, October 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Raja:2012:SDE

[170] Chandiraraj Robinson Edward Raja.

A stochastic difference equation with stationary noise on groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(5): 1075–??, October 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rosso:2012:CMR

[171] Daniele Rosso. Classic and mirabolic Robinson–Schensted–Knuth correspondence for partial flags. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(5):1090–??, October 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Seveso:2012:AFR

[172] Marco Adamo Seveso. p-adic Lfunctions and the rationality of Darmon cycles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(5):1122–??, October 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tall:2012:PMM

[173] Franklin D. Tall. PFA(S)[S]: More mutually consistent topological consequences of P F A and V = L. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(5): 1182–??, October 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aistleitner:2012:CLT

[174] Christoph Aistleitner and Christian Elsholtz. The Central Limit Theorem for subsequences in probabilistic number theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(6):1201– ??, December 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bobinski:2012:NMO

[175] Grzegorz Bobiński. Normality of maximal orbit closures for Euclidean quivers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(6):1222–??, December 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gartner:2012:DPQ

[176] Jérôme Gärtner. Darmon’s points and quaternionic Shimura varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(6): 1248–??, December 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gomes:2012:SWC

[177] Diogo Gomes and António Serra. Systems of weakly coupled Hamilton–Jacobi equations with implicit obstacles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(6): 1289–??, December 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harutyunyan:2012:UCD

[178] Ararat Harutyunyan, P. Mark Kayll, Bojan Mohar, and Liam Rafferty.

Uniquely D-colourable digraphs with large girth. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(6):1310–??, December 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Izuchi:2012:COI

[179] Kei Ji Izuchi, Quang Dieu Nguyen, and Shûichi Ohno. Composition operators induced by analytic maps to the polydisk. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(6):1329–??, December 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Killough:2012:BMH

[180] D. B. Killough and I. F. Putnam. Bowen measure from heteroclinic points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(6): 1341–??, December 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nozaki:2012:NCF

[181] Hiroshi Nozaki and Masanori Sawa.

Note on cubature formulae and designs obtained from group orbits. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(6): 1359–??, December 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Raghavan:2012:WTF

[182] Dilip Raghavan and Juris Steprans.

On weakly tight families. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(6):

1378–??, December 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rodney:2012:EWS

[183] Scott Rodney. Existence of weak solutions of linear subelliptic Dirichlet problems with rough coefficients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(6): 1395–??, December 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Selmi:2012:GWP

[184] Ridha Selmi. Global well-posedness and convergence results for 3D-regularized Boussinesq system. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(6):1415– ??, December 2012. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barto:2013:FRA

[185] Libor Barto. Finitely related algebras in congruence distributive varieties have near unanimity terms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(1):3–??, February 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blomer:2013:NVF

[186] Valentin Blomer and Farrell Brumley. Non-vanishing of L-functions, the Ramanujan conjecture, and families of Hecke characters. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(1):22–??, February 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Christensen:2013:ANC

[187] Erik Christensen, Allan M. Sinclair, Roger R. Smith, and Stuart White. C^{∗}- algebras nearly contained in type I algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(1):52–??, February 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deng:2013:FCH

[188] Shaoqiang Deng and Zhiguang Hu. On flag curvature of homogeneous Randers spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(1):66–??, February 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Felix:2013:RHG

[189] Yves Félix, Steve Halperin, and Jean-Claude Thomas. The ranks of the homotopy groups of a finite dimensional complex. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(1):82–??, February 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Francois:2013:UFR

[190] Georges François and Simon Hampe.

Universal families of rational tropical curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(1):120–??, February 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kellendonk:2013:EDD

[191] Johannes Kellendonk and Daniel Lenz.

Equicontinuous Delone dynamical systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(1):149–??, February 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lyall:2013:OPR

[192] Neil Lyall and Ákos Magyar. Optimal polynomial recurrence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(1): 171–??, February 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Penegini:2013:SAM

[193] Matteo Penegini and Francesco Polizzi. Surfaces with p_{g} = q = 2, K^{2} = 6, and Albanese map of degree 2. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(1): 195–??, February 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sauer:2013:DSU

[194] N. W. Sauer. Distance sets of Urysohn metric spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(1):222–??, February 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aguiar:2013:LTH

[195] Marcelo Aguiar and Aaron Lauve. Lagrange’s theorem for Hopf monoids in species. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(2):241–??, April 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berard:2013:ACH

[196] Vincent Bérard. Les applications conforme-harmoniques. (French) [conformalharmonic applications]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(2):266–??, April 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grafakos:2013:MFM

[197] Loukas Grafakos, Akihiko Miyachi, and Naohito Tomita. On multilinear Fourier multipliers of limited smoothness. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(2): 299–??, April 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kadets:2013:LNI

[198] Vladimir Kadets, Miguel Martín, Javier Merí, and Dirk Werner. Lushness, numerical index 1 and the daugavet property in rearrangement invariant spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65 (2):331–??, April 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Muller:2013:EPR

[199] Peter Müller and Christoph Richard. Ergodic properties of randomly coloured point sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(2):349–??, April 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

VanOrder:2013:DMC

[200] Jeanine Van Order. On the dihedral main conjectures of Iwasawa theory for Hilbert modular eigenforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(2): 403–??, April 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wilson:2013:QFC

[201] Glen Wilson and Christopher T. Woodward. Quasimap Floer cohomology for varying symplectic quotients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(2): 467–??, April 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ara:2013:CPS

[202] Pere Ara, Kenneth J. Dykema, and Mikael Rørdam. Correction of proofs in “Purely Infinite Simple C^{∗}-algebras Arising from Free Product Constructions” and a subsequent paper. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(3): 481–??, June 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bice:2013:FCA

[203] Tristan Matthew Bice. Filters in C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(3):485–??, June 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

delaCruz:2013:TVV

[204] Oscar Blasco de la Cruz and Paco Villarroya Alvarez. Transference of vector-valued multipliers on weighted L^{p}-spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(3):510–??, June 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Deitmar:2013:IIH

[205] Anton Deitmar and Ivan Horozov. Iterated integrals and higher order invariants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(3):544–??, June 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Godinho:2013:AES

[206] Leonor Godinho and M. E. Sousa-Dias. Addendum and erratum to “The Fundamental Group of S^{1}-manifolds”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65 (3):553–??, June 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Helemskii:2013:EVP

[207] A. Ya. Helemskii. Extreme version of projectivity for normed modules over sequence algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(3):559–??, June 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kallel:2013:GFG

[208] Sadok Kallel and Walid Taamallah. The geometry and fundamental group of permutation products and fat diagonals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65 (3):575–??, June 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kroo:2013:CFU

[209] A. Kroó and D. S. Lubinsky. Christoffel functions and universality in the bulk for multivariate orthogonal polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(3):600–??, June 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:2013:STD

[210] Paul W. Y. Lee. On surfaces in three dimensional contact manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(3): 621–??, June 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mezzetti:2013:LEW

[211] Emilia Mezzetti, Rosa M. Miró-Roig, and Giorgio Ottaviani. Laplace equations and the weak Lefschetz property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65 (3):634–??, June 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shemyakova:2013:PCD

[212] E. Shemyakova. Proof of the completeness of Darboux Wronskian formulae for order two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(3):655–??, June 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Strungaru:2013:BDS

[213] Nicolae Strungaru. On the Bragg diffraction spectra of a Meyer set. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(3):

675–??, June 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Taylor:2013:RSW

[214] Michael Taylor. Regularity of standing waves on Lipschitz domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(3): 702–??, June 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Adamus:2013:TCD

[215] Janusz Adamus, Serge Randriambololona, and Rasul Shafikov. Tameness of complex dimension in a real analytic set. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65 (4):721–??, August 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bernard:2013:RSD

[216] P. Bernard and M. Zavidovique. Regularization of subsolutions in discrete weak KAM theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(4):740–??, August 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Delanoe:2013:PCR

[217] Philippe Delanoë and François Rouvière.

Positively curved Riemannian locally symmetric spaces are positively squared distance curved. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(4):757–??, August 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fuller:2013:NAS

[218] Adam Hanley Fuller. Nonselfadjoint semicrossed products by Abelian semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(4):768–??, August 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Garces:2013:GTH

[219] Jorge J. Garcés and Antonio M. Peralta. Generalised triple homomorphisms and derivations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(4):783–??, August 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grandjean:2013:HLD

[220] Vincent Grandjean. On Hessian limit directions along gradient trajectories. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65 (4):808–??, August 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Guardo:2013:SPV

[221] Elena Guardo, Brian Harbourne, and Adam Van Tuyl. Symbolic powers versus regular powers of ideals of general points in P^{1} × P^{1}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(4):823–??, August 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jonsson:2013:THC

[222] Jakob Jonsson. 3-torsion in the homology of complexes of graphs of bounded degree. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(4):843–??, August 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Josuat-Verges:2013:CSL

[223] Matthieu Josuat-Vergès. Cumulants of the q-semicircular law, Tutte polynomials, and heaps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(4):863–??, August 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kawabe:2013:SHM

[224] Hiroko Kawabe. A space of harmonic maps from the sphere into the complex projective space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(4):879–??, August 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thompson:2013:EMT

[225] Alan Thompson. Explicit models for threefolds fibred by K3 surfaces of degree two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(4):905–??, August 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wang:2013:IMS

[226] Liping Wang and Chunyi Zhao. Infinitely many solutions for the prescribed boundary mean curvature problem in B^{N}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(4):927–??, August 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aholt:2013:HSC

[227] Chris Aholt, Bernd Sturmfels, and Rekha Thomas. A Hilbert scheme in computer vision. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(5):961–??, October 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chu:2013:ACH

[228] C-H. Chu and M. V. Velasco. Automatic continuity of homomorphisms in non-associative Banach algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(5): 989–??, October 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Forrest:2013:UCF

[229] Brian Forrest and Tianxuan Miao. Uniformly continuous functionals and M-weakly amenable groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(5): 1005–??, October 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goulden:2013:MHN

[230] I. P. Goulden, Mathieu Guay-Paquet, and Jonathan Novak. Monotone Hurwitz numbers in genus zero. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(5): 1020–??, October 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hu:2013:CTC

[231] Zhiguo Hu, Matthias Neufang, and Zhong-Jin Ruan. Convolution of trace class operators over locally compact quantum groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(5):1043–??, October 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kalantar:2013:QGG

[232] Mehrdad Kalantar and Matthias Neufang. From quantum groups to groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65 (5):1073–??, October 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sambou:2013:RPS

[233] Diomba Sambou. Résonances près de seuils d’opérateurs magnétiques de Pauli et de Dirac. (French) [resonances near the thresholds of magnetic operators of Pauli and Dirac]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(5):1095–??, October 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vandenbergen:2013:GSS

[234] Nicolas Vandenbergen. On the global structure of special cycles on unitary Shimura varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(5):1125–??, October 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vitagliano:2013:PDH

[235] Luca Vitagliano. Partial differential Hamiltonian systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(5):1164–??, October 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cho:2013:ASA

[236] Peter J. Cho and Henry H. Kim. Application of the strong Artin conjecture to the class number problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(6): 1201–??, December 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cruz:2013:BEC

[237] Victor Cruz, Joan Mateu, and Joan Orobitg. Beltrami equation with coefficient in Sobolev and Besov spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(6): 1217–??, December 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DeBernardi:2013:HCP

[238] Carlo Alberto De Bernardi. Higher connectedness properties of support points and functionals of convex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(6): 1236–??, December 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Iglesias-Zemmour:2013:VID

[239] Patrick Iglesias-Zemmour. Variations of integrals in diffeology. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(6):1255– ??, December 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reihani:2013:TFT

[240] Kamran Reihani. K-theory of Furstenberg transformation group C^{∗}-algebras.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 65 (6):1287–??, December 2013. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Taniguchi:2013:OFS

[241] Takashi Taniguchi and Frank Thorne.

Orbital L-functions for the space of binary cubic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(6):1320– ??, December 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wright:2013:EHD

[242] Paul Wright. Estimates of Hausdorff dimension for non-wandering sets of higher dimensional open billiards. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(6): 1384–??, December 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhao:2013:UVC

[243] Wei Zhao and Yibing Shen. A universal volume comparison theorem for Finsler manifolds and related results. Canadian Journal of Mathematics = Journal canadien de mathématiques, 65(6): 1401–??, December 2013. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Abdesselam:2014:HC

[244] Abdelmalek Abdesselam and Jaydeep Chipalkatti. On Hilbert covariants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66 (1):3–??, February 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bailey:2014:SFG

[245] Michael Bailey. Symplectic foliations and generalized complex structures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66 (1):31–??, February 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bezuglyi:2014:POF

[246] S. Bezuglyi, J. Kwiatkowski, and R. Yassawi. Perfect orderings on finite rank Bratteli diagrams. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(1):57–??, February 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Birth:2014:CCT

[247] Lidia Birth and Helge Glöckner. Continuity of convolution of test functions on Lie groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(1):102–??, February 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Caillat-Gibert:2014:ETF

[248] Shanti Caillat-Gibert and Daniel Matignon. Existence of taut foliations on Seifert fibered homology 3-spheres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66 (1):141–??, February 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Guitart:2014:MAV

[249] Xavier Guitart and Jordi Quer. Modular Abelian varieties over number fields.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 66 (1):170–??, February 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harris:2014:HDS

[250] Adam Harris and Martin Kolár. On hyperbolicity of domains with strictly pseudoconvex ends. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(1):197–??, February 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Iovanov:2014:GFA

[251] Miodrag Cristian Iovanov. Generalized Frobenius algebras and Hopf algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66 (1):205–??, February 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Broussous:2014:TDP

[252] P. Broussous. Transfert du pseudocoefficient de Kottwitz et formules de caractère pour la série discrète de GL(N) sur un corps local. (French) [Transfer of Kottwitz’s pseudo-coefficient and character formulars for the discrete series of GL(N) on a local body]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(2):241–??, April 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Eikrem:2014:RHF

[253] Kjersti Solberg Eikrem. Random harmonic functions in growth spaces and Bloch-type spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(2):284–??, April 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elekes:2014:HNS

[254] Márton Elekes and Juris Steprans. Haar null sets and the consistent reflection of non-meagreness. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(2):303–??, April 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hohlweg:2014:ABR

[255] Christophe Hohlweg, Jean-Philippe Labbé, and Vivien Ripoll. Asymptotical behaviour of roots of infinite Coxeter groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(2):323–??, April 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kellerhals:2014:MGR

[256] Ruth Kellerhals and Alexander Kolpakov. The minimal growth rate of cocompact Coxeter groups in hyperbolic 3-space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(2):354–??, April 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kim:2014:UCB

[257] Sun Kwang Kim and Han Ju Lee. Uniform convexity and Bishop–Phelps– Bollobás property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(2):373–??, April 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mashreghi:2014:CIF

[258] J. Mashreghi and M. Shabankhah. Composition of inner functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(2): 387–??, April 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mendonca:2014:US

[259] Bruno Mendonça and Ruy Tojeiro. Umbilical submanifolds of S^{n} × R. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(2): 400–??, April 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rivera-Noriega:2014:PSI

[260] Jorge Rivera-Noriega. Perturbation and solvability of initial L^{p} Dirichlet problems for parabolic equations over noncylindrical domains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(2):429–??, April 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vaz:2014:RBA

[261] Pedro Vaz and Emmanuel Wagner. A remark on BMW algebra, q-Schur algebras and categorification. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(2):453–??, April 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aguiar:2014:HPH

[262] Marcelo Aguiar and Swapneel Mahajan. On the Hadamard product of Hopf monoids. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(3):481–??, June 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arapura:2014:HTC

[263] Donu Arapura. Hodge theory of cyclic covers branched over a union of hyperplanes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(3):505–??, June 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berg:2014:LSH

[264] Chris Berg, Nantel Bergeron, Franco Saliola, Luis Serrano, and Mike Zabrocki. A lift of the Schur and Hall– Littlewood bases to non-commutative symmetric functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(3):525–??, June 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Choiy:2014:TPM

[265] Kwangho Choiy. Transfer of Plancherel measures for unitary supercuspidal representations between p-adic inner forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66 (3):566–??, June 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Eilers:2014:OTF

[266] Søren Eilers, Gunnar Restorff, and Efren Ruiz. The ordered K-theory of a full extension. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(3):596–??, June 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Giambruno:2014:CMV

[267] Antonio Giambruno, Daniela La Mattina, and Mikhail Zaicev. Classifying the minimal varieties of polynomial growth. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66 (3):625–??, June 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grigoryan:2014:HKG

[268] Alexander Grigor’yan and Jiaxin Hu.

Heat kernels and Green functions on metric measure spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(3):641–??, June 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

He:2014:IRT

[269] Jianxun He and Jinsen Xiao. Inversion of the Radon transform on the free nilpotent Lie group of step two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66 (3):700–??, June 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Durand-Cartagena:2014:WTC

[270] E. Durand-Cartagena, L. Ihnatsyeva, R. Korte, and M. Szumańska. On Whitney-type characterization of approximate differentiability on metric measure spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(4):721–??, August 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hrusak:2014:NCD

[271] Michael Hrusák and Jan van Mill.

Nearly countable dense homogeneous spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(4):743–??, August 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See addendum [272].

Hrusak:2014:AT

[272] Michael Hrusák and Jan van Mill. Addendum to ‘Nearly Countable Dense Homogeneous Spaces’. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(4):759–??, August 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [271].

Hu:2014:RKP

[273] Shengda Hu and Manuele Santoprete. Regularization of the Kepler problem on the three-sphere. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(4):760–??, August 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Izmestiev:2014:IRC

[274] Ivan Izmestiev. Infinitesimal rigidity of convex polyhedra through the second derivative of the Hilbert–Einstein functional. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(4):783–??, August 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kim:2014:SSG

[275] Byoung Du Kim. Signed-Selmer groups over the Z_{p}^{2}-extension of an imaginary quadratic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(4):826–??, August 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kuo:2014:MVT

[276] Wentang Kuo, Yu-Ru Liu, and Xiaomei Zhao. Multidimensional Vinogradovtype estimates in function fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(4): 844–??, August 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levandovskyy:2014:QDH

[277] Viktor Levandovskyy and Anne V. Shepler. Quantum Drinfeld Hecke algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(4):874–??, August 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See corrigendum [278].

Levandovskyy:2014:CEI

[278] Viktor Levandovskyy and Anne V. Shepler. Corrigendum to example in “Quantum Drinfeld Hecke Algebras”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(4): 902–??, August 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [277].

Sargsyan:2014:NTM

[279] Grigor Sargsyan and Nam Trang. Nontame mice from tame failures of the unique branch hypothesis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(4): 903–??, August 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stankewicz:2014:TSC

[280] James Stankewicz. Twists of Shimura curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(4):924–??, August 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baird:2014:MSV

[281] Thomas Baird. Moduli spaces of vector bundles over a real curve: Z/2– Betti numbers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(5):961–??, October 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beuzart-Plessis:2014:EFE

[282] Raphaël Beuzart-Plessis. Expression d’un facteur epsilon de paire par une formule intégrale. (French) [Expression of a pair-epsilon factor by an integral formula]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(5):993–??, October 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Holmes:2014:RWD

[283] Mark Holmes and Thomas S. Salisbury. Random walks in degenerate random environments. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(5):1050–??, October 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lanphier:2014:VTT

[284] Dominic Lanphier and Howard Skogman. Values of twisted tensor Lfunctions of automorphic forms over imaginary quadratic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(5): 1078–??, October 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Li:2014:DED

[285] Dong Li, Guixiang Xu, and Xiaoyi Zhang. On the dispersive estimate for the Dirichlet Schrödinger propagator and applications to energy critical NLS. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(5):1110–??, October 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Plevnik:2014:MPC

[286] Lucijan Plevnik and Peter Semrl. Maps preserving complementarity of closed subspaces of a Hilbert space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(5): 1143–??, October 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rotger:2014:GRF

[287] Victor Rotger and Carlos de Vera-Piquero. Galois representations over fields of moduli and rational points on Shimura curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(5):1167–??, October 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Adler:2014:LRF

[288] Jeffrey D. Adler and Joshua M. Lansky. Lifting representations of finite reductive groups I: Semisimple conjugacy classes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(6):1201–??, December 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benitez:2014:MGD

[289] Teresa Cortadellas Benítez and Carlos D’Andrea. Minimal generators of the defining ideal of the Rees algebra associated with a rational plane parametrization with μ = 2. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(6): 1225–??, December 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feigin:2014:SDF

[290] Evgeny Feigin, Michael Finkelberg, and Peter Littelmann. Symplectic degenerate flag varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(6):1250– ??, December 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Henniart:2014:TC

[291] Guy Henniart and Vincent Sécherre. Types et contragrédientes. (French) [Types and contragredients]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(6): 1287–??, December 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koskivirta:2014:CRS

[292] Jean-Stefan Koskivirta. Congruence relations for Shimura varieties associated with GU(n − 1, 1). Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(6): 1305–??, December 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mohar:2014:OCT

[293] Bojan Mohar and Petr Skoda. Obstructions of connectivity two for embedding graphs into the torus. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(6): 1327–??, December 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Osekowski:2014:SLI

[294] Adam Osekowski. Sharp localized inequalities for Fourier multipliers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(6): 1358–??, December 2014. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wu:2014:WCM

[295] Xinfeng Wu. Weighted Carleson measure spaces associated with different homogeneities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(6):1382–??, December 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhang:2014:GKE

[296] Xi Zhang and Xiangwen Zhang. Generalized Kähler–Einstein metrics and energy functionals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 66(6):1413–??, December 2014. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Alfonseca:2015:LCI

[297] M. Angeles Alfonseca and Jaegil Kim.

On the local convexity of intersection bodies of revolution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(1):3–??, February 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Asadollahi:2015:BDC

[298] Javad Asadollahi, Rasool Hafezi, and Razieh Vahed. Bounded derived categories of infinite quivers: Grothendieck duality, reflection functor. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(1): 28–??, February 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barron:2015:VLA

[299] Tatyana Barron, Dmitry Kerner, and Marina Tvalavadze. On varieties of Lie algebras of maximal class. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(1): 55–??, February 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bousch:2015:PDC

[300] Thierry Bousch. Une propriété de domination convexe pour les orbites sturmiennes. (French) [A property of convex domination for Sturmian orbits]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(1): 90–??, February 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chang:2015:WPN

[301] Jui-En Chang and Ling Xiao. The Weyl problem with nonnegative Gauss curvature in hyperbolic space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(1): 107–??, February 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clouatre:2015:UES

[302] Raphaël Clouâtre. Unitary equivalence and similarity to Jordan models for weak contractions of class C_{0}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(1): 132–??, February 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lescop:2015:HIC

[303] Christine Lescop. On homotopy invariants of combings of three-manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67 (1):152–??, February 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McReynolds:2015:GSC

[304] D. B. McReynolds. Geometric spectra and commensurability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(1): 184–??, February 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Murty:2015:TCA

[305] V. Kumar Murty and Vijay M.

Patankar. Tate cycles on Abelian varieties with complex multiplication. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(1): 198–??, February 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Szpruch:2015:SGS

[306] Dani Szpruch. Symmetric genuine spherical Whittaker functions on \overline{GSp}_{2n}(F). Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(1):214–??, February 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Agler:2015:GHF

[307] Jim Agler and John E. McCarthy. Global holomorphic functions in several noncommuting variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(2):??, April 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bell:2015:SML

[308] Jason P. Bell and Jeffrey C. Lagarias. A Skolem–Mahler–Lech theorem for iterated automorphisms of K-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67 (2):??, April 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bellaiche:2015:UEI

[309] Joël Bellaïche. Unitary eigenvarieties at isobaric points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(2):??, April 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bernardes:2015:HDG

[310] Nilson C. Bernardes, Jr. and Rômulo M. Vermersch. Hyperspace dynamics of generic maps of the Cantor space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67 (2):??, April 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Colombo:2015:MOT

[311] Maria Colombo, Luigi De Pascale, and Simone Di Marino. Multimarginal optimal transport maps for one-dimensional repulsive costs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(2):??, April 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Graham:2015:FPF

[312] Robert Graham and Mikael Pichot. A free product formula for the sofic dimension. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(2):??, April 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hua:2015:RAE

[313] Jiajie Hua and Huaxin Lin. Rotation algebras and the Exel trace formula. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(2):??, April 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Samart:2015:MML

[314] Detchat Samart. Mahler measures as linear combinations of L-values of multiple modular forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(2):??, April 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Santoprete:2015:MSP

[315] Manuele Santoprete, Jürgen Scheurle, and Sebastian Walcher. Motion in a symmetric potential on the hyperbolic plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(2):??, April 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

anHuef:2015:ACT

[316] Astrid an Huef and Robert John Archbold. The C∗-algebras of compact transformation groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(3):??, June 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:2015:LPP

[317] Peter Borwein, Stephen Choi, Ron Ferguson, and Jonas Jankauskas. On Littlewood polynomials with prescribed number of zeros inside the unit disk. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(3):??, June 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brugalle:2015:OAT

[318] Erwan Brugallé and Kristin Shaw. Obstructions to approximating tropical curves in surfaces via intersection theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(3):??, June 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2015:TVO

[319] Fulin Chen, Yun Gao, Naihuan Jing, and Shaobin Tan. Twisted vertex operators and unitary Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(3):??, June 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Drappeau:2015:SFE

[320] Sary Drappeau. Sommes friables d’exponentielles et applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(3): ??, June 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gonzalez:2015:PAC

[321] Jose Luis Gonzalez and Kalle Karu. Projectivity in algebraic cobordism. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(3): ??, June 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lim:2015:GSG

[322] Meng Fai Lim and V. Kumar Murty.

Growth of Selmer groups of CM Abelian varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(3):??, June 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nishinou:2015:TDT

[323] Takeo Nishinou. Toric degenerations, tropical curve, and Gromov–Witten invariants of Fano manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(3):??, June 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhang:2015:GIG

[324] Tong Zhang. Geography of irregular Gorenstein 3-folds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(3):??, June 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Allen:2015:DCH

[325] Peter Allen, Julia Böttcher, Jan Hladký, and Diana Piguet. A density Corrádi-Hajnal theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(4):721–??, August 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carey:2015:SFN

[326] Alan L. Carey, Victor Gayral, John Phillips, Adam Rennie, and Fedor Sukochev. Spectral flow for nonunital spectral triples. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(4):759–??, August 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nasso:2015:SCE

[327] Mauro Di Nasso, Isaac Goldbring, Renling Jin, Steven Leth, Martino Lupini, and Karl Mahlburg. On a sumset conjecture of Erdős. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(4):795–??, August 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elliott:2015:AIE

[328] George A. Elliott and Zhuang Niu. All irrational extended rotation algebras are AF algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(4):810–??, August 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaniuth:2015:BSE

[329] Eberhard Kaniuth. The Bochner– Schoenberg–Eberlein property and spectral synthesis for certain Banach algebra products. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(4):827–??, August 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kock:2015:FAR

[330] Bernhard Köck and Joseph Tait.

Faithfulness of actions on Riemann– Roch spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(4):848–??, August 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lin:2015:MDS

[331] Huaxin Lin. Minimal dynamical systems on connected odd dimensional spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67 (4):870–??, August 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mok:2015:OFS

[332] Chung Pang Mok and Fucheng Tan.

Overconvergent families of Siegel– Hilbert modular forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(4):893–??, August 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pan:2015:CMJ

[333] Ivan Edgardo Pan and Aron Simis. Cremona maps of de Jonquières type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(4): 923–??, August 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roth:2015:PMP

[334] Oliver Roth. Pontryagin’s maximum principle for the Loewner equation in higher dimensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(4):942–??, August 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Abuaf:2015:OBS

[335] Roland Abuaf and Ada Boralevi. Orthogonal bundles and skew-Hamiltonian matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(5):961–??, October 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Amini:2015:CBD

[336] Massoud Amini, George A. Elliott, and Nasser Golestani. The category of Bratteli diagrams. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(5):990–??, October 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ashraf:2015:RSP

[337] Samia Ashraf, Haniya Azam, and Barbu Berceanu. Representation stability of power sets and square free polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67 (5):1024–??, October 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dubickas:2015:EFC

[338] Arturas Dubickas, Min Sha, and Igor Shparlinski. Explicit form of Cassels p-adic embedding theorem for number fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(5):1046–??, October 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ducrot:2015:FTC

[339] Arnaud Ducrot, Pierre Magal, and Ousmane Seydi. A finite-time condition for exponential trichotomy in infinite dynamical systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(5):1065–??, October 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mine:2015:MCC

[340] Kotaro Mine and Atsushi Yamashita.

Metric compactifications and coarse structures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(5):1091–??, October 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nohara:2015:GSB

[341] Yuichi Nohara and Kazushi Ueda. Goldman systems and bending systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67 (5):1109–??, October 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nystedt:2015:OPA

[342] Patrik Nystedt and Johan Öinert. Outer partial actions and partial skew group rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(5):1144–??, October 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhang:2015:NTM

[343] Junqiang Zhang, Jun Cao, Renjin Jiang, and Dachun Yang. Non-tangential maximal function characterizations of Hardy spaces associated with degenerate elliptic operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(5):1161–??, October 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aluffi:2015:CCS

[344] Paolo Aluffi and Eleonore Faber. Chern classes of splayed intersections. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(6): 1201–??, December 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Balwe:2015:AMM

[345] Chetan Balwe. p-adic and Motivic measure on Artin n-stacks. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(6): 1219–??, December 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barros:2015:LSA

[346] Carlos Braga Barros, Victor Rocha, and Josiney Souza. Lyapunov stability and attraction under equivariant maps.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 67 (6):1247–??, December 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carcamo:2015:SES

[347] Cristian Carcamo and Claudio Vidal.

Stability of equilibrium solutions in planar Hamiltonian difference systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67 (6):1270–??, December 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Charlesworth:2015:TFF

[348] Ian Charlesworth, Brent Nelson, and Paul Skoufranis. On two-faced families of non-commutative random variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(6):1290–??, December 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cojocaru:2015:DFE

[349] Alina Carmen Cojocaru and Andrew Michael Shulman. The distribution of the first elementary divisor of the reductions of a generic Drinfeld module of arbitrary rank. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(6): 1326–??, December 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Trillos:2015:RCE

[350] Nicolas Garcia Trillos and Dejan Slepcev. On the rate of convergence of empirical measures in ∞- transportation distance. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(6):1358–??, December 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Graczyk:2015:SLS

[351] Piotr Graczyk, Todd Kemp, and Jean-Jacques Loeb. Strong logarithmic Sobolev inequalities for logsubharmonic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(6):1384–??, December 2015. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kawakami:2015:FTP

[352] Yu Kawakami. Function-theoretic properties for the Gauss maps of various classes of surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 67(6):1411– ??, December 2015. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boden:2016:SCI

[353] Hans Ulysses Boden and Cynthia L. Curtis. The SL(2, C) Casson invariant for knots and the Â-polynomial. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(1): 3–??, February 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bonfanti:2016:ASA

[354] Matteo Alfonso Bonfanti and Bert van Geemen. Abelian surfaces with an automorphism and quaternionic multiplication. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(1):24–??, February 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Breton:2016:SSU

[355] David J. Fernández Bretón. Strongly summable ultrafilters, union ultrafilters, and the trivial sums property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(1): 44–??, February 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ishida:2016:LBE

[356] Hirotaka Ishida. A lower bound on the Euler–Poincaré characteristic of certain surfaces of general type with a linear pencil of hyperelliptic curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(1): 67–??, February 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jaffe:2016:PPD

[357] Ethan Y. Jaffe. Pathological phenomena in Denjoy–Carleman classes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(1): 88–??, February 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kopotun:2016:CAJ

[358] Kirill Kopotun, Dany Leviatan, and Igor Shevchuk. Constrained approximation with Jacobi weights. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(1):109–??, February 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shiozawa:2016:LER

[359] Yuichi Shiozawa. Lower escape rate of symmetric jump-diffusion processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68 (1):129–??, February 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stavrova:2016:NSF

[360] Anastasia Stavrova. Non-stable K_{1}- functors of multiloop groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(1): 150–??, February 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Takeda:2016:MTP

[361] Shuichiro Takeda. Metaplectic tensor products for automorphic representation of G̃L(r). Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(1):179–??, February 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Allermann:2016:RET

[362] Lars Allermann, Simon Hampe, and Johannes Rau. On rational equivalence in tropical geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(2):241–??, April 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Calixto:2016:EMQ

[363] Lucas Calixto, Adriano Moura, and Alistair Savage. Equivariant map queer Lie superalgebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(2):258–??, April 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

daSilva:2016:ADP

[364] Genival da Silva, Jr., Matt Kerr, and Gregory Pearlstein. Arithmetic of degenerating principal variations of Hodge structure: Examples arising from mirror symmetry and middle convolution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68 (2):280–??, April 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Daws:2016:CAQ

[365] Matthew Daws. Categorical aspects of quantum groups: Multipliers and intrinsic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(2):309–??, April 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Demchenko:2016:KCF

[366] Oleg Demchenko and Alexander Gurevich. Kernels in the category of formal group laws. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(2):334–??, April 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fite:2016:FDQ

[367] Francesc Fité, Josep González, and Joan Carles Lario. Frobenius distribution for quotients of Fermat curves of prime exponent. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(2):361–??, April 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Garibaldi:2016:BQF

[368] Skip Garibaldi and Daniel K. Nakano. Bilinear and quadratic forms on rational modules of split reductive groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68 (2):395–??, April 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kohen:2016:HPC

[369] Daniel Kohen and Ariel Pacetti. Heegner points on Cartan non-split curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68 (2):422–??, April 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martins:2016:GIC

[370] Luciana de Fátima Martins and Kentaro Saji. Geometric invariants of cuspidal edges. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(2):445–??, April 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sadykov:2016:WPM

[371] Rustam Sadykov. The weak b-principle: Mumford conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(2):463–??, April 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bacher:2016:NRI

[372] Roland Bacher and Christophe Reutenauer. Number of right ideals and a q-analogue of indecomposable permutations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68 (3):481–??, June 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Biswas:2016:IST

[373] Indranil Biswas, Tomás L. Gómez, and Marina Logares. Integrable systems and Torelli theorems for the moduli spaces of parabolic bundles and parabolic Higgs bundles. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(3):504–??, June 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Emamizadeh:2016:ORS

[374] Behrouz Emamizadeh, Amin Farjudian, and Mohsen Zivari-Rezapour. Optimization related to some nonlocal problems of Kirchhoff type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(3):521–??, June 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Garcia-Armas:2016:SIC

[375] Mario Garcia-Armas. Strongly incompressible curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(3):541–??, June 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gras:2016:RLN

[376] Georges Gras. Les θ-régulateurs locaux d’un nombre algébrique : Conjectures p-adiques. (French) [The local θ regulators of an algebraic number: p-adic conjectures]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(3):571–??, June 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ingram:2016:RHB

[377] Patrick Ingram. Rigidity and height bounds for certain post-critically finite endomorphisms of P^{N}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(3):625–??, June 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Klartag:2016:DCA

[378] Bo’az Klartag, Gady Kozma, Peter Ralli, and Prasad Tetali. Discrete curvature and Abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(3): 655–??, June 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martinez-de-la-Vega:2016:MCD

[379] Veronica Martínez de-la Vega and Christopher Mouron. Monotone classes of dendrites. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(3):675–??, June 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Skalski:2016:QFI

[380] Adam Skalski and Piotr Soltan. Quantum families of invertible maps and related problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(3):698–??, June 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chandee:2016:FEC

[381] Vorrapan Chandee, Chantal David, Dimitris Koukoulopoulos, and Ethan Smith. The frequency of elliptic curve groups over prime finite fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(4): 721–??, August 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Colesanti:2016:LRP

[382] Andrea Colesanti, Eugenia Saorín Gómez, and Jesus Yepes Nicolás. On a linear refinement of the Prékopa– Leindler inequality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(4):762–??, August 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Doran:2016:TDL

[383] Charles F. Doran and Andrew Harder.

Toric degenerations and Laurent polynomials related to Givental’s Landau– Ginzburg models. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(4):784–??, August 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Guo:2016:CSI

[384] Xiaoli Guo and Guoen Hu. On the commutators of singular integral operators with rough convolution kernels. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(4): 816–??, August 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gupta:2016:CAC

[385] Sanjiv Kumar Gupta and Kathryn Hare.

Characterizing the absolute continuity of the convolution of orbital measures in a classical Lie algebra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(4):841–??, August 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ostrovskii:2016:MSA

[386] Mikhail Ostrovskii and Beata Randrianantoanina. Metric spaces admitting low-distortion embeddings into all n-dimensional Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(4): 876–??, August 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sugiyama:2016:EHC

[387] Shingo Sugiyama and Masao Tsuzuki. Existence of Hilbert cusp forms with non-vanishing L-values. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(4):908–??, August 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Greenberg:2016:AFC

[388] Matthew Greenberg and Marco Seveso. p-adic families of cohomological modular forms for indefinite quaternion algebras and the Jacquet–Langlands correspondence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(5):961–??, October 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Izumi:2016:Q

[389] Masaki Izumi, Scott Morrison, and David Penneys. Quotients of A_{2} ∗ T_{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68 (5):999–??, October 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Phillips:2016:CVI

[390] John Phillips and Iain Raeburn. Centrevalued index for Toeplitz operators with noncommuting symbols. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(5): 1023–??, October 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Runde:2016:PDL

[391] Volker Runde and Ami Viselter. On positive definiteness over locally compact quantum groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(5):1067–??, October 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:2016:SM

[392] Benjamin H. Smith. Singular Gmonopoles on S^{1} ×Σ. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(5):1096–??, October 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stange:2016:IPE

[393] Katherine E. Stange. Integral points on elliptic curves and explicit valuations of division polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(5):1120–??, October 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yattselev:2016:SAH

[394] Maxim L. Yattselev. Strong asymptotics of Hermite–Padé approximants for Angelesco systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(5):1159–??, October 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Banks:2016:MAU

[395] Jessica Banks and Matt Rathbun. Monodromy action on unknotting tunnels in fiber surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(6):1201–??, December 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brasca:2016:ECP

[396] Riccardo Brasca. Eigenvarieties for cuspforms over PEL type Shimura varieties with dense ordinary locus. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(6): 1227–??, December 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cascante:2016:SNE

[397] Carme Cascante, Joan Fàbrega, and Joaquín M. Ortega. Sharp norm estimates for the Bergman operator from weighted mixed-norm spaces to weighted Hardy spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(6):1257–??, December 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ehrig:2016:RSF

[398] Michael Ehrig and Catharina Stroppel. 2-row Springer fibres and Khovanov diagram algebras for type D. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(6): 1285–??, December 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jiang:2016:NPM

[399] Feida Jiang, Neil S. Trudinger, and Ni Xiang. On the Neumann problem for Monge–Ampère type equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(6): 1334–??, December 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Papikian:2016:OQJ

[400] Mihran Papikian and Joseph Rabinoff.

Optimal quotients of Jacobians with toric reduction and component groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68 (6):1362–??, December 2016. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zydor:2016:VIF

[401] Michal Zydor. La variante infinitésimale de la formule des traces de Jacquet– Rallis pour les groupes unitaires. Canadian Journal of Mathematics = Journal canadien de mathématiques, 68(6): 1382–??, December 2016. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ghahramani:2017:BIB

[402] F. Ghahramani and S. Zadeh. Bipositive isomorphisms between Beurling algebras and between their second dual algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(1):3–??, February 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grinberg:2017:DCO

[403] Darij Grinberg. Dual creation operators and a dendriform algebra structure on the quasisymmetric functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(1): 21–??, February 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartz:2017:IPM

[404] Michael Hartz. On the isomorphism problem for multiplier algebras of Nevanlinna–Pick spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(1): 54–??, February 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kamgarpour:2017:NCL

[405] Masoud Kamgarpour. On the notion of conductor in the local geometric Langlands correspondence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(1): 107–??, February 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levin:2017:NAC

[406] Aaron Levin and Julie Tzu-Yueh Wang. On non-Archimedean curves omitting few components and their arithmetic analogues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(1):130–??, February 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levinson:2017:ODS

[407] Jake Levinson. One-dimensional Schubert problems with respect to osculating flags. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(1):143–??, February 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pan:2017:FLT

[408] Shu-Yen Pan. L-functoriality for local theta correspondence of supercuspidal representations with unipotent reduction. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(1):186–??, February 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zheng:2017:CRF

[409] Tao Zheng. The Chern–Ricci flow on Oeljeklaus–Toma manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(1): 220–??, February 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Adamus:2017:FDS

[410] Janusz Adamus and Hadi Seyedinejad. Finite determinacy and stability of flatness of analytic mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69 (2):241–??, April 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brandes:2017:SAE

[411] Julia Brandes and Scott T. Parsell.

Simultaneous additive equations: Repeated and differing degrees. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(2): 258–??, April 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2017:CPS

[412] Xianghong Chen and Andreas Seeger. Convolution powers of Salem measures with applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(2):284–??, April 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DeBernardi:2017:TNS

[413] Carlo Alberto De Bernardi and Libor Veselý. Tilings of normed spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69 (2):321–??, April 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Garbagnati:2017:KSQ

[414] Alice Garbagnati. On K3 surface quotients of K3 or Abelian surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(2): 338–??, April 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaftal:2017:SCP

[415] Victor Kaftal, Ping Wong Ng, and Shuang Zhang. Strict comparison of positive elements in multiplier algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69 (2):373–??, April 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Klep:2017:FFT

[416] Igor Klep and Spela Spenko. Free function theory through matrix invariants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(2):408–??, April 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:2017:NDC

[417] Hun Hee Lee and Sang gyun Youn.

New deformations of convolution algebras and Fourier algebras on locally compact groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(2):434–??, April 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marquis:2017:ITH

[418] Timothée Marquis and Karl-Hermann Neeb. Isomorphisms of twisted Hilbert loop algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(2):453–??, April 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cordero-Erausquin:2017:TIL

[419] Dario Cordero-Erausquin. Transport inequalities for log-concave measures, quantitative forms and applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69 (3):481–??, June 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fischer:2017:SBA

[420] Vera Fischer and Diego Alejandro Mejia.

Splitting, bounding, and almost disjointness can be quite different. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(3): 502–??, June 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ganguly:2017:DTF

[421] Arijit Ganguly and Anish Ghosh. Dirichlet’s theorem in function fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69 (3):532–??, June 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartglass:2017:FPC

[422] Michael Hartglass. Free product C^{∗}- algebras associated with graphs, free differentials, and laws of loops. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(3): 548–??, June 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lee:2017:RIF

[423] Jungyun Lee and Yoonjin Lee. Regulators of an infinite family of the simplest quartic function fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(3):579–??, June 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mauduit:2017:DS

[424] Christian Mauduit, Joël Rivat, and András Sárközy. On the digits of sumsets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(3):595–??, June 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moon:2017:MPS

[425] Han-Bom Moon. Mori’s program for \overline{M}_{0,7} with symmetric divisors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(3): 613–??, June 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Oikhberg:2017:ADP

[426] Timur Oikhberg and Pedro Tradacete.

Almost disjointness preservers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(3): 650–??, June 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ovchinnikov:2017:TCS

[427] Alexey Ovchinnikov and Michael Wibmer. Tannakian categories with semigroup actions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(3):687–??, June 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Allison:2017:WIK

[428] Bruce Allison, John Faulkner, and Oleg Smirnov. Weyl images of Kantor pairs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69 (4):721–??, August 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Choi:2017:WOT

[429] Suyoung Choi and Hanchul Park. Wedge operations and torus symmetries II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69 (4):767–??, August 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Diacu:2017:CBP

[430] Florin Diacu. The classical N-body problem in the context of curved space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69 (4):790–??, August 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gunther:2017:NFR

[431] Christian Günther and Kai-Uwe Schmidt. L^{q} norms of Fekete and related polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(4):807–??, August 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lei:2017:AGB

[432] Antonio Lei, David Loeffler, and Sarah Livia Zerbes. On the asymptotic growth of Bloch–Kato–Shafarevich–Tate groups of modular forms over cyclotomic extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(4):826–??, August 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pronk:2017:ETG

[433] Dorette Pronk and Laura Scull. Erratum: Translation Groupoids and Orbifold Cohomology. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(4):851–??, August 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [34].

Saanouni:2017:GNG

[434] Tarek Saanouni. Global and non global solutions for some fractional heat equations with pure power nonlinearity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69 (4):854–??, August 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Xiao:2017:ASC

[435] Jie Xiao and Deping Ye. Anisotropic Sobolev capacity with fractional order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69 (4):873–??, August 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Xu:2017:MPA

[436] Bin Xu. On Moeglin’s parametrization of Arthur packets for p-adic quasisplit Sp(N) and SO(N). Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(4):890–??, August 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Andrade:2017:DRK

[437] Jaime Andrade, Nestor Dávila, Ernesto Pérez-Chavela, and Claudio Vidal. Dynamics and regularization of the Kepler problem on surfaces of constant curvature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(5):961–??, October 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bremner:2017:CRP

[438] Murray Bremner and Vladimir Dotsenko. Classification of regular parametrized one-relation operads. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69 (5):992–??, October 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlen:2017:SBM

[439] Eric Carlen and Francesco Maggi. Stability for the Brunn–Minkowski and Riesz rearrangement inequalities, with applications to Gaussian concentration and finite range non-local isoperimetry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69 (5):1036–??, October 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Crann:2017:ACI

[440] Jason Crann. Amenability and covariant injectivity of locally compact quantum groups II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(5):1064–??, October 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jiang:2017:ACW

[441] Yin Jiang. Absolute continuity of Wasserstein barycenters over Alexandrov spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(5):1087–??, October 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ng:2017:CCH

[442] P. W. Ng and P. Skoufranis. Closed convex hulls of unitary orbits in certain simple real rank zero C^{∗}-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(5): 1109–??, October 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sikiric:2017:SDP

[443] Mathieu Dutour Sikirić. The seven dimensional perfect Delaunay polytopes and Delaunay simplices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(5):1143–??, October 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Varma:2017:RIO

[444] Sandeep Varma. On residues of intertwining operators in cases with prehomogeneous nilradical. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(5):1169–??, October 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Abe:2017:CPL

[445] Tetsuya Abe and Keiji Tagami. Characterization of positive links and the s-invariant for links. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(6):1201– ??, December 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Almeida:2017:AHL

[446] Víctor Almeida, Jorge J. Betancor, and Lourdes Rodríguez-Mesa. Anisotropic Hardy–Lorentz spaces with variable exponents. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(6):1219–??, December 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Favacchio:2017:MFR

[447] Giuseppe Favacchio and Elena Guardo.

The minimal free resolution of fat almost complete intersections in P^{1}×P^{1}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(6): 1274–??, December 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Folha:2017:WTS

[448] Abigail Folha and Carlos Peñafiel. Weingarten type surfaces in H^{2} ×R and S^{2} × R. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69 (6):1292–??, December 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fricain:2017:AOS

[449] Emmanuel Fricain and Rishika Rupam. On asymptotically orthonormal sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(6):1312–??, December 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Harrison-Trainor:2017:CFE

[450] Matthew Harrison-Trainor, Alexander Melnikov, and Russell Miller. On computable field embeddings and difference closed fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(6):1338– ??, December 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nikolidakis:2017:ESB

[451] Eleftherios Nikolaos Nikolidakis. Extremal sequences for the Bellman function of the dyadic maximal operator and applications to the Hardy operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(6):1364–??, December 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pasnicu:2017:WIP

[452] Cornel Pasnicu and N. Christopher Phillips. The weak ideal property and topological dimension zero. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69(6): 1385–??, December 2017. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Semrl:2017:OSP

[453] Peter Semrl. Order and spectrum preserving maps on positive operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 69 (6):1422–??, December 2017. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Benaych-Georges:2018:FME

[454] Florent Benaych-Georges, Guillaume Cébron, and Jean Rochet. Fluctuation of matrix entries and application to outliers of elliptic matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(1):3–??, February 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bosa:2018:CPC

[455] Joan Bosa and Henning Petzka. Comparison properties of the Cuntz semigroup and applications to C∗-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70 (1):26–??, February 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dantas:2018:BPB

[456] Sheldon Dantas, Domingo García, Manuel Maestre, and Miguel Martín. The Bishop–Phelps–Bollobás property for compact operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(1):53–??, February 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dow:2018:NVP

[457] Alan Dow and Franklin D. Tall. Normality versus paracompactness in locally compact spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(1):74–??, February 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Farashahi:2018:CAL

[458] Arash Ghaani Farashahi. A class of abstract linear representations for convolution function algebras over homogeneous spaces of compact groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(1): 97–??, February 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ha:2018:SPS

[459] Junsoo Ha. Smooth polynomial solutions to a ternary additive equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70 (1):117–??, February 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hajir:2018:IFC

[460] Farshid Hajir and Christian Maire. On the invariant factors of class groups in towers of number fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(1): 142–??, February 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hakl:2018:PSI

[461] Robert Hakl and Manuel Zamora. Periodic solutions of an indefinite singular equation arising from the Kepler problem on the sphere. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(1):173–??, February 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kitchloo:2018:C

[462] Nitu Kitchloo, Vitaly Lorman, and W. Stephen Wilson. The ER(2)- cohomology of BZ/(2^{q}) and CP^{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70 (1):191–??, February 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Speissegger:2018:QIA

[463] Patrick Speissegger. Quasianalytic Ilyashenko algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(1):218–??, February 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bocherer:2018:WMK

[464] Siegfried Böcherer, Toshiyuki Kikuta, and Sho Takemori. Weights of the mod p kernel of the theta operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70 (2):241–??, April 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bulens:2018:RMC

[465] Hector Cordova Bulens, Pascal Lambrechts, and Don Stanley. Rational models of the complement of a subpolyhedron in a manifold with boundary. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(2):265–??, April 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Eilers:2018:GCG

[466] Søren Eilers, Gunnar Restorff, Efren Ruiz, and Adam P. W. Sørensen. Geometric classification of graph C∗- algebras over finite graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(2): 294–??, April 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Manon:2018:TGF

[467] Christopher Manon. Toric geometry of SL_{2}(C) free group character varieties from outer space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(2):354–??, April 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Osaka:2018:JSA

[468] Hiroyuki Osaka and Tamotsu Teruya. The Jiang–Su absorption for inclusions of unital C∗-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(2):400–??, April 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See erratum [566].

Perez-Chavela:2018:ETR

[469] Ernesto Pérez-Chavela and Juan Manuel Sánchez-Cerritos. Euler-type relative equilibria in spaces of constant curvature and their stability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(2):426–??, April 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zhang:2018:EOS

[470] Chao Zhang. Ekedahl–Oort strata for good reductions of Shimura varieties of Hodge type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(2):451–??, April 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Asakura:2018:CPC

[471] Masanori Asakura and Noriyuki Otsubo.

CM periods, CM regulators and hypergeometric functions, I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(3):481–??, June 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chen:2018:BTG

[472] Yanni Chen, Don Hadwin, Zhe Liu, and Eric Nordgren. A Beurling theorem for generalized Hardy spaces on a multiply connected domain. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(3):515–??, June 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ciesielski:2018:FPT

[473] Krzysztof Chris Ciesielski and Jakub Jasinski. Fixed point theorems for maps with local and pointwise contraction properties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(3):538–??, June 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohen:2018:TRO

[474] Jonathan Cohen. Transfer of representations and orbital integrals for inner forms of GL_{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(3):595–??, June 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Luo:2018:SLL

[475] Ye Luo and Madhusudan Manjunath. Smoothing of limit linear series of rank one on saturated metrized complexes of algebraic curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(3):628–??, June 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Matringe:2018:GFR

[476] Nadir Matringe and Omer Offen. Gamma factors, root numbers, and distinction. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(3):683–??, June 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Xia:2018:ARC

[477] Eugene Z. Xia. The algebraic de Rham cohomology of representation varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70 (3):702–??, June 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bao:2018:DSC

[478] Guanlong Bao, Nihat Gökhan Gögüs, and Stamatis Pouliasis. On Dirichlet spaces with a class of superharmonic weights. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(4):721–??, August 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bijakowski:2018:PHI

[479] Stephane Bijakowski. Partial Hasse invariants, partial degrees, and the canonical subgroup. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(4):742–??, August 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Du:2018:MFC

[480] Jie Du and Zhonghua Zhao. Multiplication formulas and canonical bases for quantum affine gl_{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(4):773–??, August 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Giannopoulos:2018:ISA

[481] Apostolos Giannopoulos, Alexander Koldobsky, and Petros Valettas. Inequalities for the surface area of projections of convex bodies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(4):804–??, August 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hare:2018:LDM

[482] Kathryn Hare, Kevin Hare, and Michael Ka Shing Ng. Local dimensions of measures of finite type II — measures without full support and with nonregular probabilities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(4):824–??, August 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ivorra:2018:NMC

[483] Florian Ivorra and Takao Yamazaki. Nori motives of curves with modulus and Laumon 1-motives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(4):868–??, August 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Luo:2018:SFL

[484] Caihua Luo. Spherical fundamental lemma for metaplectic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(4): 898–??, August 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McDiarmid:2018:EMG

[485] Colin McDiarmid and David R. Wood. Edge-maximal graphs on surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70 (4):925–??, August 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yuan:2018:CFN

[486] Rirong Yuan. On a class of fully nonlinear elliptic equations containing gradient terms on compact Hermitian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(4):943–??, August 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ara:2018:UNC

[487] Pere Ara and Joan Claramunt. Uniqueness of the von Neumann continuous factor. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(5):961–??, October 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Conway:2018:ECB

[488] Anthony Conway. An explicit computation of the Blanchfield pairing for arbitrary links. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(5):983–??, October 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elazar:2018:SFR

[489] Boaz Elazar and Ary Shaviv. Schwartz functions on real algebraic varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70 (5):1008–??, October 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elduque:2018:OEI

[490] Alberto Elduque. Order 3 elements in G_{2} and idempotents in symmetric composition algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(5):1038–??, October 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martin:2018:CMF

[491] Kimball Martin. Congruences for modular forms mod 2 and quaternionic Sideal classes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(5):1076–??, October 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mullner:2018:RSS

[492] Clemens Müllner. The Rudin–Shapiro sequence and similar sequences are normal along squares. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(5):1096–??, October 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rushworth:2018:DKH

[493] William Rushworth. Doubled Khovanov homology. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(5):1130–??, October 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Viada:2018:EMD

[494] Evelina Viada. An explicit Manin– Dem’janenko theorem in elliptic curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70 (5):1173–??, October 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bickerton:2018:FMW

[495] Robert T. Bickerton and Evgenios T. A. Kakariadis. Free multivariate w∗- semicrossed products: Reflexivity and the bicommutant property. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(6): 1201–??, December 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Clouatre:2018:UPO

[496] Raphaël Clouâtre. Unperforated pairs of operator spaces and hyperrigidity of operator systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(6):1236– ??, December 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fricain:2018:RSC

[497] Emmanuel Fricain, Andreas Hartmann, and William T. Ross. Range spaces of co-analytic Toeplitz operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(6): 1261–??, December 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Geroldinger:2018:LSL

[498] Alfred Geroldinger and Qinghai Zhong.

Long sets of lengths with maximal elasticity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(6):1284–??, December 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Macourt:2018:MES

[499] Simon Macourt, Ilya D. Shkredov, and Igor E. Shparlinski. Multiplicative energy of shifted subgroups and bounds on exponential sums with trinomials in finite fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(6):1319–??, December 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:2018:RDS

[500] Jerrod Manford Smith. Relative discrete series representations for two quotients of p-adic GL_{n}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(6):1339– ??, December 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tuxanidy:2018:NPH

[501] Aleksandr Tuxanidy and Qiang Wang. A new proof of the Hansen–Mullen irreducibility conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(6):1373– ??, December 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Xiao:2018:SFV

[502] Stanley Yao Xiao. Square-free values of decomposable forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(6): 1390–??, December 2018. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yeats:2018:SCC

[503] Karen Yeats. A special case of completion invariance for the c_{2} invariant of a graph. Canadian Journal of Mathematics = Journal canadien de mathématiques, 70(6):1416–??, December 2018. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bernardi:2019:BAP

[504] Olivier Bernardi, Nicolas Curien, and Grégory Miermont. A Boltzmann approach to percolation on random triangulations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(1):1–43, February 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/boltzmann-
approach-to-percolation-on-random-
triangulations/907258D8620557E4A95D55AD80C35B74.

Camere:2019:CYQ

[505] Chiara Camere, Alice Garbagnati, and Giovanni Mongardi. Calabi–Yau quotients of hyperkähler four-folds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (1):45–92, February 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/calabiyau-quotients-of-
hyperkahler-fourfolds/1F1952F82C0713424E18C9B93C42A696.

Courtney:2019:ECX

[506] Kristin Courtney and Tatiana Shulman. Elements of C^{∗}-algebras attaining their norm in a finite-dimensional representation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(1):93–111, February 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/elements-
of-cast-algebras-attaining-their-
norm-in-a-finitedimensional-representation/
9791B0E9815632B3C320153A931A8186.

deVerclos:2019:CSC

[507] Rémi de Joannis de Verclos, Ross J.

Kang, and Lucas Pastor. Colouring squares of claw-free graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (1):113–129, February 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/colouring-squares-of-clawfree-
graphs/CC21DF02708EBB427347374278BC6274.

Glockner:2019:CID

[508] Helge Glöckner. Completeness of infinite-dimensional Lie groups in their left uniformity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(1):131– 152, February 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/completeness-of-infinitedimensional-
lie-groups-in-their-left-uniformity/
A85E1B10990A991A1730E98BBEF2DAFA.

Knightly:2019:WDL

[509] Andrew Knightly and Caroline Reno. Weighted distribution of low-lying zeros of GL(2) L-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (1):153–182, February 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/weighted-distribution-of-
lowlying-zeros-of-gl2-l-functions/
C5DD2FE6BBFEEA872430A27BC2FD5D84.

Li:2019:BQC

[510] Hui Li and Dilian Yang. Boundary quotient C^{∗}-algebras of products of odometers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(1):183–212, February 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/boundary-quotient-textcast-
algebras-of-products-of-odometers/
A46FA034F5F788775694A62E3E037FF9.

Shimada:2019:ESA

[511] Ichiro Shimada. On an Enriques surface associated with a quartic Hessian surface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(1):213–246, February 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-an-enriques-surface-
associated-with-a-quartic-hessian-
surface/7F4ED300013922C9D14F240FEA1B5DC4.

See corrigendum [567].

Anonymous:2019:CVIa

[512] Anonymous. CJM volume 71 issue 1 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (1):f1–f2, February 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-71-issue-1-cover-
and-front-matter/DAAFDD9952B6DD5DD6CCDE7A0A92073E.

Anonymous:2019:CVIb

[513] Anonymous. CJM volume 71 issue 1 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (1):b1–b2, February 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/

                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-71-issue-1-cover-
                                        and-back-matter/0D3719AF618277850B363740C3E94774.

Bosser:2019:LPR

[514] Vincent Bosser and Éric Gaudron. Logarithmes des points rationnels des variétés abéliennes. (French) [Logarithms of Abelian rational points]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(2):247–298, April 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/logarithmes-des-points-
                                        rationnels-des-varietes-abeliennes/
                                        2B88A5E69C1547F9E49AB75D38918FC5.

Dyer:2019:WOC

[515] Matthew Dyer. On the weak order of Coxeter groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(2): 299–336, April 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/on-the-weak-order-of-coxeter-
                                        groups/5BA500A691F73B68906EBCD63AD4CFAE.

Georgescu:2019:IFS

[516] Magdalena Cecilia Georgescu. Integral formula for spectral flow for psummable operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(2):337–379, April 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

                                        URL https://www.cambridge.org/
                                        core/journals/canadian-journal-

of-mathematics/article/integral-
formula-for-spectral-flow-for-p-
summable-operators/B238C1993435E9DB83E6BD14B109AC9A.

Handelman:2019:NAT

[517] David Handelman. Nearly Approximate Transitivity (AT) for circulant matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(2):381–415, April 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/nearly-approximate-transitivity-
at-for-circulant-matrices/D70EC7F9ADD081CEE36FA5793A365364.

Karpukhin:2019:SPD

[518] Mikhail A. Karpukhin. The Steklov problem on differential forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (2):417–435, April 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/steklov-problem-on-differential-
forms/CAD648C54499E5A02A527ACF403EDC25.

Lambie-Hanson:2019:FAD

[519] Chris Lambie-Hanson and Assaf Rinot.

A forcing axiom deciding the generalized Souslin hypothesis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(2):437–470, April 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/forcing-
axiom-deciding-the-generalized-
souslin-hypothesis/282980A496B9C1911B578C7118AD68EC.

Wang:2019:ASA

[520] Zhenjian Wang. On algebraic surfaces associated with line arrangements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(2):471–499, April 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/on-algebraic-surfaces-associated-
                                        with-line-arrangements/05D45DB58D6F7D839478149FC213BDAE.

Anonymous:2019:CVIc

[521] Anonymous. CJM volume 71 issue 2 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (2):f1–f2, April 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-71-issue-2-cover-
                                        and-front-matter/000B55D7E3994A39A76EB52286B3CB58.

Anonymous:2019:CVId

[522] Anonymous. CJM volume 71 issue 2 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (2):b1–b2, April 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-71-issue-2-cover-
                                        and-back-matter/FF90E03DB6D06AD933C372F498AD768F.

Astashkin:2019:ISC

[523] Sergey V. Astashkin, Karol Lesnik, and Lech Maligranda. Isomorphic structure of Cesàro and Tandori spaces.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(3):501–532, June 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/isomorphic-structure-of-
cesaro-and-tandori-spaces/5F412C2A5AFB88497CC1B0B3F6C67A55.

Cohen:2019:LCS

[524] David Bruce Cohen. Lipschitz 1-connectedness for some solvable Lie groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(3):533–555, June 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/lipschitz-
1connectedness-for-some-solvable-
lie-groups/BA519D462E2FB4F02C590905421D452D.

Galetto:2019:DRS

[525] Federico Galetto, Anthony Vito Geramita, and David Louis Wehlau. Degrees of regular sequences with a symmetric group action. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(3):557–578, June 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/degrees-
of-regular-sequences-with-a-symmetric-
group-action/81AB19192D5AAD7680B02A0853F8F3BB.

Green:2019:MSX

[526] Ben Joseph Green and Sofia Lindqvist.

Monochromatic solutions to x + y = z^{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(3):579–605, June 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/monochromatic-solutions-
                                        to-xyz2/FA809E6B6EDBC5BE02F2930AAD556406.

Han:2019:MML

[527] Yanchang Han, Yongsheng Han, Ji Li, and Chaoqiang Tan. Marcinkiewicz multipliers and Lipschitz spaces on Heisenberg groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(3):607– 627, June 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/marcinkiewicz-multipliers-
                                        and-lipschitz-spaces-on-heisenberg-
                                        groups/9C656415E4E0992502D9C24FAB2E695F.

He:2019:SLL

[528] Xiang He. Smoothing of limit linear series on curves and metrized complexes of pseudocompact type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(3):629–658, June 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

                                        URL https://www.cambridge.org/
                                        core/journals/canadian-journal-
                                        of-mathematics/article/smoothing-
                                        of-limit-linear-series-on-curves-
                                        and-metrized-complexes-of-pseudocompact-
                                        type/6CCE79CF518371506257CE2092C332CD.

Mingo:2019:FPT

[529] James A. Mingo and Mihai Popa. Freeness and the partial transposes of Wishart random matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(3): 659–681, June 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/freeness-and-the-partial-
transposes-of-wishart-random-matrices/
9C807D7530735A92D822D97BFB46393C.

Scaduto:2019:MTC

[530] Christopher W. Scaduto and Matthew Stoffregen. The mod two cohomology of the moduli space of rank two stable bundles on a surface and skew Schur polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(3):683–715, June 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/mod-two-
cohomology-of-the-moduli-space-
of-rank-two-stable-bundles-on-a-
surface-and-skew-schur-polynomials/
E0C26E79F0D4DF2EB1A5DFF858C95416.

Stokke:2019:FSC

[531] Ross Stokke. Fourier spaces and completely isometric representations of Arens product algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(3):717–747, June 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/fourier-
spaces-and-completely-isometric-
representations-of-arens-product-
algebras/4C9E61012E3EB3ECD00E1731BB74467C.

Anonymous:2019:CVIe

[532] Anonymous. CJM volume 71 issue 3 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (3):f1–f2, June 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-71-issue-3-cover-
                                        and-front-matter/3F4D9B80CFF4735D2E571A7298DD1A26.

Anonymous:2019:CVIf

[533] Anonymous. CJM volume 71 issue 3 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (3):b1–b2, June 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cjm-volume-71-issue-3-cover-
                                        and-back-matter/5FCEE81006EBCB3B94C1A51BF6EF9193.

Bourhim:2019:LMP

[534] Abdellatif Bourhim and Constantin Costara. Linear maps preserving matrices of local spectral radius zero at a fixed vector. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(4):749– 771, August 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/linear-maps-preserving-
                                        matrices-of-local-spectral-radius-
                                        zero-at-a-fixed-vector/9A59DD2C1466ADCD8FC2598ED6516F14.

Cahn:2019:POR

[535] Jordan Cahn, Rafe Jones, and Jacob Spear. Powers in orbits of rational functions: Cases of an arithmetic dynamical Mordell–Lang conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(4):773–817, August 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL https://www.cambridge.org/
core/journals/canadian-journal-
of-mathematics/article/powers-in-
orbits-of-rational-functions-cases-
of-an-arithmetic-dynamical-mordelllang-
conjecture/C294E2DF514470392D5A466A03B6D469.

Kaygorodov:2019:VTD

[536] Ivan Kaygorodov and Yury Volkov.

The variety of two-dimensional algebras over an algebraically closed field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (4):819–842, August 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/variety-of-twodimensional-
algebras-over-an-algebraically-
closed-field/BBCF5D27C25551F8CA86F1A8BEC6369B.

Kuribayashi:2019:BVA

[537] Katsuhiko Kuribayashi and Luc Menichi.

The Batalin–Vilkovisky algebra in the string topology of classifying spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (4):843–889, August 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/

                                        article/batalinvilkovisky-algebra-
                                        in-the-string-topology-of-classifying-
                                        spaces/99AA5CC400B20221A71E30F51DD9C5CD.

Mihara:2019:CAC

[538] Tomoki Mihara. Cohomological approach to class field theory in arithmetic topology. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(4):891– 935, August 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cohomological-approach-
                                        to-class-field-theory-in-arithmetic-
                                        topology/06AFE3BA42FE082831E381DB08FDDFB2.

Phan:2019:LEW

[539] Tuoc Phan. Lorentz estimates for weak solutions of quasi-linear parabolic equations with singular divergencefree drifts. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(4):937–982, August 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

                                        URL https://www.cambridge.org/
                                        core/journals/canadian-journal-
                                        of-mathematics/article/lorentz-
                                        estimates-for-weak-solutions-of-
                                        quasilinear-parabolic-equations-
                                        with-singular-divergencefree-drifts/
                                        5719EBF8DFEA01C340B7E33E72BD83C7.

Wang:2019:PCS

[540] Xing Wang and Chunjie Zhang. Pointwise convergence of solutions to the Schrödinger equation on manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (4):983–995, August 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/pointwise-convergence-of-
solutions-to-the-schrodinger-equation-
on-manifolds/F24B0C567B3E32CA6308418A43914316.

Anonymous:2019:CVIg

[541] Anonymous. CJM volume 71 issue 4 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(4): f1–f2, August 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-71-issue-4-cover-
and-front-matter/65C3E1E7259F3FA12D3DDDBD4891FDD2.

Anonymous:2019:CVIh

[542] Anonymous. CJM volume 71 issue 4 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (4):b1–b2, August 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-71-issue-4-cover-
and-back-matter/FA9B0C2874F3BE17BC1EF1A6EB4D73DD.

Arzhantseva:2019:GIP

[543] Goulnara Arzhantseva and Cornelia Drutu. Geometry of infinitely presented small cancellation groups and quasi-homomorphisms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(5):997– 1018, October 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/geometry-of-infinitely-
                                        presented-small-cancellation-groups-
                                        and-quasihomomorphisms/DB70B68118C9CCA2E76F118356107796.

Salazar:2019:AF

[544] Daniel Barrera Salazar and Chris Williams. p-adic L-functions for GL_{2}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (5):1019–1059, October 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/p-adic-l-functions-for-
                                        textgl2/8B2CEFE6D536BB75523A2F6471193285.

Brundan:2019:BTD

[545] Jonathan Brundan, Jonathan Comes, and Jonathan Robert Kujawa. A basis theorem for the degenerate affine oriented Brauer–Clifford supercategory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (5):1061–1101, October 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/basis-theorem-for-the-degenerate-
                                        affine-oriented-brauerclifford-
                                        supercategory/5A69C40B569D8AD84D835007C504309F.

Cameron:2019:GCR

[546] Jan Cameron and Roger R. Smith. A Galois correspondence for reduced crossed products of simple C^{∗}-algebras by discrete groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(5):1103– 1125, October 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/galois-correspondence-for-
reduced-crossed-products-of-simple-
textcast-algebras-by-discrete-groups/
EE5EC67CFB2D19038582AB6903E93502.

See corrigendum [565].

Gurevich:2019:PSR

[547] Nadya Gurevich and Avner Segal.

Poles of the standard L-function of G_{2} and the Rallis–Schiffmann lift. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(5): 1127–1161, October 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/poles-of-the-standard-mathcall-
function-of-g2-and-the-rallisschiffmann-
lift/6361BD7D53EDFE5A41954D39A114CF53.

Hartl:2019:LSD

[548] Urs Hartl and Rajneesh Kumar Singh.

Local shtukas and divisible local Anderson modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(5):1163– 1207, October 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/local-shtukas-and-divisible-
local-anderson-modules/CCCD6EA24B89FF3C6322F5A6A72391D9.

Iacono:2019:DPM

[549] Donatella Iacono and Marco Manetti.

On deformations of pairs (manifold, coherent sheaf). Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(5):1209– 1241, October 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-deformations-of-pairs-
manifold-coherent-sheaf/B017822BED1B8D6202816C2E10C0A30D.

Matsumoto:2019:ACO

[550] Kengo Matsumoto. Asymptotic continuous orbit equivalence of Smale spaces and Ruelle algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(5): 1243–1296, October 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/asymptotic-continuous-orbit-
equivalence-of-smale-spaces-and-
ruelle-algebras/80775F8A86A5E3018CDAB51F048A0881.

Anonymous:2019:CVIi

[551] Anonymous. CJM volume 71 issue 5 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(5): f1–f2, October 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-71-issue-5-cover-
and-front-matter/3BDEC8006C97750C9B07BC2860B234BA.

Anonymous:2019:CVIj

[552] Anonymous. CJM volume 71 issue 5 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (5):b1–b2, October 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-71-issue-5-cover-
and-back-matter/D4AE92CB6A9DEA9029B3AC3E7E0AE9C9.

Barlow:2019:GUS

[553] Martin T. Barlow and Antal A. Járai.

Geometry of uniform spanning forest components in high dimensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(6): 1297–1321, December 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/geometry-of-uniform-spanning-
forest-components-in-high-dimensions/
E9BE6FB7AFA0D166CFCAF60B696B2DEC.

Bary-Soroker:2019:CTP

[554] Lior Bary-Soroker and Jakob Stix. Cubic twin prime polynomials are counted by a modular form. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(6):1323– 1350, December 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cubic-twin-prime-polynomials-
are-counted-by-a-modular-form/30221F68B7F55D76C3CF0DF41E48B3D7.

Bump:2019:CBI

[555] Daniel Bump and Maki Nakasuji. Casselman’s basis of Iwahori vectors and Kazhdan–Lusztig polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(6): 1351–1366, December 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/casselmans-basis-of-iwahori-
                                        vectors-and-kazhdanlusztig-polynomials/
                                        47E6822DD0D7458A1F9EF659555AC3A0.

Chang:2019:CAY

[556] Der-Chen Chang, Shu-Cheng Chang, Yingbo Han, and Jingzhi Tie. A CR analogue of Yau’s conjecture on pseudoharmonic functions of polynomial growth. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (6):1367–1394, December 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/cr-analogue-of-yaus-conjecture-
                                        on-pseudoharmonic-functions-of-
                                        polynomial-growth/41E2F7EF0755C530F6EC232BD83C39A9.

Chapdelaine:2019:AIC

[557] Hugo Chapdelaine and Radan Kucera. Annihilators of the ideal class group of a cyclic extension of an imaginary quadratic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(6):1395– 1419, December 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/annihilators-of-the-ideal-
                                        class-group-of-a-cyclic-extension-
                                        of-an-imaginary-quadratic-field/
                                        C12CD26099B095A3A08E58AE8E3BE032.

Dantas:2019:PBP

[558] Sheldon Dantas, Vladimir Kadets, Sun Kwang Kim, Han Ju Lee, and Miguel Martín. On the pointwise Bishop–Phelps–Bollobás property for operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(6):1421– 1443, December 2019. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/on-the-pointwise-bishopphelpsbollobas-
property-for-operators/A54148CCB84EFA34173D9E798BC8AC10.

Dyachenko:2019:UCT

[559] Mikhail Dyachenko, Askhat Mukanov, and Sergey Tikhonov. Uniform convergence of trigonometric series with general monotone coefficients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(6): 1445–1463, December 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/uniform-convergence-of-
trigonometric-series-with-general-
monotone-coefficients/715E1E8331CDCC909E8E80BF4874B8B0.

Furuya:2019:TMA

[560] Jun Furuya, T. Makoto Minamide, and Yoshio Tanigawa. Titchmarsh’s method for the approximate functional equations for ζ′(s)^{2}, ζ(s)ζ′′(s), and ζ′(s)ζ′′(s). Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (6):1465–1493, December 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/titchmarshs-method-for-
the-approximate-functional-equations-
for-unicodestixx1d701prime-s2-unicodestixx1d701sunicodestixx1d701prime-

                                        prime-s-and-unicodestixx1d701prime-
                                        sunicodestixx1d701prime-prime-s/
                                        7A973F187A082F9AC4C7AAC4048A76C8.

Liu:2019:FPS

[561] Ricky Ini Liu, Alejandro H. Morales, and Karola Mészáros. Flow polytopes and the space of diagonal harmonics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71(6): 1495–1521, December 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/flow-polytopes-and-the-
                                        space-of-diagonal-harmonics/53927DF1CF0A47E2D85D1A44EFAC73E2.

Mackaaij:2019:TCS

[562] Marco Mackaaij and Daniel Tubbenhauer. Two-color Soergel calculus and simple transitive 2-representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (6):1523–1566, December 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/twocolor-soergel-calculus-
                                        and-simple-transitive-2representations/
                                        9911E3B037D3C3CA1942CB09C525BAE8.

Anonymous:2019:CVIk

[563] Anonymous. CJM volume 71 issue 6 cover and front matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (6):f1–f2, December 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/

article/cjm-volume-71-issue-6-cover-
and-front-matter/A2EFBA5065C9C992DAC8BB215C458D31.

Anonymous:2019:CVIl

[564] Anonymous. CJM volume 71 issue 6 cover and back matter. Canadian Journal of Mathematics = Journal canadien de mathématiques, 71 (6):b1–b2, December 2019. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/cjm-volume-71-issue-6-cover-
and-back-matter/0E5C310A54FFA80A8DE02E164B8A1137.

Cameron:2020:CGC

[565] Jan Cameron and Roger R. Smith. Corrigendum to: A Galois Correspondence for Reduced Crossed Products of Simple C^{∗}-algebras by Discrete Groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 72(2):557–562, April 2020. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

/www.cambridge.org/core/journals/
canadian-journal-of-mathematics/
article/corrigendum-to-a-galois-
correspondence-for-reduced-crossed-
products-of-simple-textcast-algebras-
by-discrete-groups/C9B9CAAF1F3BE9677AF9595EE3DD5CC7.

See [546].

Osaka:2021:EJA

[566] Hiroyuki Osaka and Tamotsu Teruya.

Erratum: The Jiang–Su Absorption for Inclusions of Unital C∗-algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 73 (1):293–295, February 2021. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/erratum-the-jiangsu-absorption-
                                        for-inclusions-of-unital-calgebras/
                                        5818143B89D6DF74DD853FD7C9E0075A.

See [468].

Shimada:2022:CES

[567] Ichiro Shimada. Corrigendum: On an Enriques surface associated with a quartic Hessian surface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 74 (2):603–605, April 2022. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL https:/

                                        /www.cambridge.org/core/journals/
                                        canadian-journal-of-mathematics/
                                        article/corrigendum-on-an-enriques-
                                        surface-associated-with-a-quartic-
                                        hessian-surface/88812D1442FA6F51029DAD8BF5BCFBCC.

See [511].